### v1

# The Generalization Ridge: Information Flow in Natural Language Generation

Abstract
Transformer-based language models have achieved state-of-the-art performance in natural language generation (NLG) tasks, yet their internal mechanisms for synthesizing task-relevant information remain insufficiently understood. While prior studies suggest that intermediate layers often yield more generalizable representations than final layers, how this generalization ability emerges and propagates across layers during training remains unclear. To address this gap, we propose InfoRidge, an information-theoretic framework, to characterize how predictive information—the mutual information between hidden representations and target outputs—varies across depth. Estimating this quantity enables us to trace the flow of task-relevant information throughout the model during training. Our experiments across various models and datasets reveal a consistent non-monotonic trend: predictive information peaks in upper-middle layers—forming a generalization ridge—before declining in final layers, reflecting a transition between generalization and memorization. To further investigate this phenomenon, we introduce residual scaling coefficients—trainable scalar parameters applied to each residual block—which serve as functional probes for assessing the relative importance of individual transformer layers. These coefficients reveal that, under distribution shift, models downweight final layers and increasingly rely on ridge layers, highlighting their role in generalization. Together, these findings offer new insights into the internal mechanisms of transformers and underscore the critical role of intermediate layers in supporting generalization.
1 Introduction
Transformer-based language models have achieved remarkable performance in natural language generation (NLG) tasks such as summarization, machine translation, and dialogue generation vaswani2017attention ; dong2022survey . Nevertheless, we lack a rigorous understanding of how these models acquire and synthesize task-relevant information during training.
A growing body of research has shown that intermediate layers in deep neural networks often surpass final layers in terms of representational quality and generalization performance liu2019linguistic ; voita2019bottom ; ansuini2019intrinsic ; ahrens2023read ; uselis2025intermediate . In language models, intermediate layers often encode richer semantic and more robust features than final layers fan2024not ; jin2024exploring ; skean2025layer . However, questions still remain: In NLG, how does information evolve across layers during training, and how are different layers of the network functionally organized to support generalization versus memorization?
To investigate these questions, we propose InfoRidge, an information-theoretic framework, to analyze information flow in language models. Building on matrix-based mutual information estimation giraldo2014measures , our approach quantifies how predictive signals transform across layers during training. We center our analysis on two complementary quantities. The first is the predictive information at each layer, defined as the mutual information $I(Z_{\ell};Y)$ between the hidden representation $Z_{\ell}$ at layer $\ell$ and the next-token label $Y$, which reflects how much task-relevant information is preserved. The second is the incremental information gain, denoted as $I(\Delta Z_{\ell};Y)$, where $\Delta Z_{\ell}$ is the residual changes between successive layer $\ell$ and $\ell-1$. This measures the additional predictive information introduced by each transformer block.
Using these tools, we uncover a non-monotonic trend: predictive information rises through early and middle layers, peaks in the upper-middle, and then declines in later layers. We term this phenomenon the generalization ridge, indicating a structural division of labor across depth. Intermediate layers appear to specialize in encoding abstract, generalizable features, while later layers increasingly reorient representations toward task-specific alignment and memorization. By further analyzing incremental information gain, we find that certain middle layers introduce the largest increases in predictive information, marking them as key contributors to the emergence of the ridge. Such layer-wise specialization offers insight into how generalization and memorization are distributed within the network.
To further validate this interpretation, we introduce residual scaling coefficients—learnable scalar parameters $\beta_{\ell}$ applied to each residual block—while keeping all other model weights fixed, drawing inspiration from prior work on layer-wise adaptation liu2019self ; menghani2024laurel . A higher $\beta_{\ell}$ value indicates greater reliance on the corresponding layer’s output during prediction. These coefficients act as functional probes, revealing how the model redistributes layer-wise importance under different data distribution. Under in-distribution training, deeper layers retain higher residual weights, reflecting the model’s reliance on memorized, task-specific features. When evaluated under distribution shift, models reduce reliance on late layers and increase reliance on the ridge, further supporting its role in generalization.
To understand the formation of the generalization ridge, we analyze both attention patterns and model capacity. Attention flow analysis shows that ridge layers attend more directly to semantically informative tokens, suggesting that the ridge is not only a statistical peak but also functionally aligned with meaningful computation. In contrast, deeper layers increasingly focus on positional tokens, consistent with a shift toward memorization. Beyond functional behavior, we also investigate the conditions under which the ridge emerges. Our depth ablation results show that the ridge only emerges beyond a certain depth threshold. Below this threshold, predictive information increases monotonically, indicating that sufficient capacity is a prerequisite for generalization ridge to emerge.
Contributions.
Our work provides a unified perspective on how predictive information is structured across depth in transformer-based language models for natural language generation tasks:
  1. 1.
    We propose InfoRidge, an information-theoretic framework, to quantify both predictive information and incremental information gain across layers, offering a dynamic view of how task-relevant signals emerge and evolve during training.
  2. 2.
    Through this framework, we identify the generalization ridge—a mid-layer peak in predictive information—revealing a transition between generalization and memorization.
  3. 3.
    We propose residual scaling coefficients as functional probes for layer importance, uncovering a shift in reliance from final layers to intermediate layers under distribution shift, consistent with the ridge’s role in supporting generalization.
These findings suggest that transformer depth reflects functional stratification: intermediate layers encode core abstractions for generalization, while final layers increasingly tailor representations for memorization—often at the cost of robustness. This insight informs the design of more interpretable and adaptable architectures under distribution shift.
2 Related Work
Understanding how information is encoded and transformed across layers in deep models has been widely studied through probing classifiers alain2016understanding , attention flow analysis vig2019analyzing , and information-theoretic tools such as the information bottleneck shwartz2017opening , mutual information estimation goldfeld2019estimating , and matrix-based entropy methods giraldo2014measures . These methods provide various lenses to quantify representational capacity, abstraction, and invariance across layers.
A growing body of research has shown that intermediate layers in deep networks often outperform final layers in terms of representational quality and task performance ansuini2019intrinsic ; yosinski2014transferable ; uselis2025intermediate ; ahrens2023read ; ando2023use . In language models, mid-depth layers tend to capture richer semantic or robust features than output layers liu2019linguistic ; voita2019bottom ; jin2024exploring ; fan2024not . These findings challenge the assumption that deeper layers always yield better representations, revealing instead a peak in representational utility at intermediate depths. This pattern holds across settings such as transfer learning yosinski2014transferable , continual learning ahrens2023read , and out-of-distribution generalization uselis2025intermediate . Recent work by Skean et al. skean2025layer proposes a unified framework that evaluates representation quality using entropy, curvature, and invariance. Collectively, these results highlight the importance of intermediate layers for abstraction and generalization, motivating our analysis of layer-wise information dynamics through an information-theoretic lens. Differently, we focus on information flow for next token generation, specifically examining mutual information between the last input token and the target label. Our goal is to understand how generalization and memorization dynamics evolve during training from information-theoretic perspective.
3 InfoRidge: Information Estimation Framework
Figure 1: Overview of InfoRidge. (1) Extract internal representations at each layer and compute residual changes between successive layers. (2) Extract the target token embedding. (3) Estimate predictive information and incremental information gain.
We propose InfoRidge, an information-theoretic framework that uses mutual information to quantify how predictive information propagates through transformers layers during training in NLG.
Motivating Insight.
Prior work has shown that internal representations in deep neural networks tend to align most closely with the true data distribution at an intermediate layer he2024information . By employing the Wasserstein distance villani2008optimal , this alignment is shown to reach a minimum at a specific depth—referred to as the generalization funnel layer. At this point, the Min Wasserstein Generalization Bound he2024information ensures that the upper bound on the generalization gap—defined as the expected difference between the population and empirical risks—is minimized. This highlights the critical role of intermediate layers in supporting generalization.
Hypothesis.
Building on this insight, we hypothesize that there exists a specific intermediate transformer layer that encodes the most generalizable representations for next-token prediction, characterized by maximal mutual information with the target label.
Hypothesis: Generalization Ridge.  There exists an intermediate layer $\ell^{*}\in\{1,\dots,L\}$ such that the mutual information between the hidden state and the target label peaks at that layer:
$\ell^{*}=\arg\max_{\ell}I(Z_{\ell};Y).$
We refer to this layer as the generalization ridge. This ridge layer aligns most strongly with generalizable features and serve as robust predictors under distribution shift, whereas later layers increasingly specialize in memorization.
InfoRidge Overview.
To empirically investigate this hypothesis, we propose InfoRidge, an information estimation framework, that characterizes how predictive information evolves across transformer layers. Specifically, we estimate two key quantities:
  • Predictive Information $I(Z_{\ell};Y)$: the mutual information between the hidden state at layer $\ell$ and the target token. This quantity measures how much information about the true next token is contained in the layer’s full representation. A high value indicates that the layer encodes a strong and direct signal relevant to the prediction task.
  • Incremental Information Gain $I(\Delta Z_{\ell};Y)$: the information introduced by the residual transformation at layer $\ell$, where $\Delta Z_{\ell}=Z_{\ell}-Z_{\ell-1}$. This captures the additional predictive signal gained through the residual transformation at layer $\ell$, isolating how much new task-relevant information is introduced on top of the previous layer’s representation.
Together, these metrics allow us to track both the accumulation and transformation of task-relevant information throughout the network.
Notation and Setup.
We consider a transformer model with $L$ residual blocks. Given an input sequence $x_{1:T}$ of length $T$, $z^{(\ell)}_{i}\in\mathbb{R}^{d}$ denotes the hidden state at the last token position of the $i$th input sequence in layer $\ell$, for $\ell=1,\dots,L$. For next-token prediction, the ground-truth label is denoted by $y_{i}\in\mathbb{R}^{d}$, corresponding to the embedding of the true next token from the vocabulary $\mathcal{V}$. The residual transformation introduced at layer $\ell$ is defined as $\delta z_{i}^{(\ell)}=z_{i}^{(\ell)}-z_{i}^{(\ell-1)}$.
Across a batch of $N$ sequences, we collect the representations:
$Z_{\ell}=\{z^{(\ell)}_{i}\}_{i=1}^{N},\quad Y=\{y_{i}\}_{i=1}^{N},\quad\Delta Z_{\ell}=\{\delta z_{i}^{(\ell)}\}_{i=1}^{N},$
where all vectors are $\ell_{2}$-normalized.
Computational Flow Overview.
Figure 1 illustrates the workflow used to extract intermediate representations for information analysis. In the first forward pass, a batch of $N$ input sequences $x_{1:T}$ is fed into the transformer to obtain hidden states at each layer. From these, we extract the final-token representations $Z_{\ell}$ and compute the residual changes $\Delta Z_{\ell}$ by differencing consecutive layer outputs. In the second forward pass, each input is concatenated with its ground-truth next token $y$, and we extract the corresponding label embedding $Y$ from the output of the embedding layer. These representations are then used to compute two information-theoretic quantities: the Predictive Information $I(Z_{\ell};Y)$, and the Incremental Information Gain $I(\Delta Z_{\ell};Y)$. We estimate both $I(Z_{\ell};Y)$ and $I(\Delta Z_{\ell};Y)$ using Equation 1 and 2, detailed below.
Matrix-Based Mutual Information.
We apply the matrix-based framework giraldo2014measures to estimate mutual information. Let $\mathcal{U}$ be a random variable, from which we draw a set of vectors $U=\{\mathbf{u}_{i}\}_{i=1}^{N}\subset\mathbb{R}^{d}$. A positive-definite Gram matrix $G_{U}\in\mathbb{R}^{N\times N}$ is computed using a Gaussian kernel $\kappa$ with bandwidth set to 1 and the matrix is then trace-normalized to satisfy $\operatorname{tr}(G_{U})=1$. The matrix-based Rényi entropy (with order $\alpha=1$) is then given by:
$H(\mathcal{U})\approx H(U)=-\operatorname{tr}(G_{U}\log G_{U}).$ (1)
Specifically, $G_{U}$ is constructed with entries $(G_{U})_{ij}=\exp\left(-\frac{\|u_{i}-u_{j}\|^{2}}{2\sigma^{2}}\right)$ and then trace-normalized.
Let $G_{\mathcal{U}}$ and $G_{\mathcal{V}}$ be the trace-normalized Gram matrices for two random variables $\mathcal{U}$ and $\mathcal{V}$, respectively. The mutual information between them is computed as:
$I(\mathcal{U};\mathcal{V})\approx I(G_{\mathcal{U}};G_{\mathcal{V}})=H(G_{\mathcal{U}})+H(G_{\mathcal{V}})-H(G_{\mathcal{U}}\circ G_{\mathcal{V}}),$ (2)
where $\circ$” denotes the Hadamard (elementwise) product. Mathematical details are in Appendix A.
4 Experimental Setup
Models.
We evaluate four models: GPT–2 Small (117M) radford2019language , GPT–2 Medium (345M) radford2019language , Qwen–2.5 0.5B qwen2 , and LLaMA 3.1 8B meta_llama_3.1_8b . All models are fine-tuned on NLG tasks. Each model shares weights between the input token embedding and the output language modeling head.
Datasets.
We assess model behavior across three tasks reformulated as NLG problems: CLUTRR (sinha2019clutrr, ), a relational reasoning benchmark; ECQA (aggarwal2021explanations, ), a commonsense QA benchmark; and Synthetic Arithmetic, a controlled dataset designed to disentangle task-relevant signal from noise. Dataset statistics and implementation details are provided in Appendix B and C.
[…]
5 Generalization Ridge: Layer-wise Mutual Information Training Dynamics
Understanding how layers encode task-relevant information is key to uncovering the internal mechanisms that support generalization in deep language models. In this section, we trace the evolution of two complementary forms of mutual information—Predictive Information and Incremental Information Gain—across transformer depth and training time, revealing a consistent structure in information flow and highlighting the generalization–memorization trade-off.
5.1 Predictive Information: Information Peaks at Intermediate Layers
We investigate how predictive information—defined as the mutual information between hidden representations and target labels—evolves across transformer layers. Specifically, we calculate the matrix–based mutual information $I(Z_{\ell};Y)$ at each transformer layer $\ell$, capturing the amount of label-relevant information encoded during forward propagation.
(a) GPT-2 Small on Synthetic
(b) GPT-2 Medium on CLUTRR
(c) Qwen-2.5-0.5B on ECQA
(d) LLaMA-3.1-8B on ECQA
Figure 2: Evolution of predictive information $I(Z_{\ell};Y)$, with lighter curves indicating later epochs. Each curve exhibits a three-phase trend: early layers rise, mid layers peak, and late layers decline.
Figure 2 tracks the trajectory of the predictive information $I(Z_{\ell};Y)$ between the hidden representation at depth $\ell$ and the target label $Y$ throughout training, while Table 1 reports the downstream accuracy obtained when we early exit after a given layer. [^1] The in distribution split is generated with $K_{\text{id}}{=}13$; out of distribution splits use a uniformly–sampled $K_{\text{ood}}\in\{5,\dots,25\}\setminus\{13\}$. Additional results are presented in Appendix E.
Three-phase information dynamics.
Across models and tasks, predictive information curves exhibit a consistent three-phase pattern:
Progressive Accrual (early layers). In the initial layers, $I(Z_{\ell};Y)$ gradually increases, corresponding to basic feature extraction without substantial task-level comprehension. This aligns with the near-zero accuracy observed in Table 1 for these layers.
Table 1: Information Dynamics and Layer-wise Performance (GPT-2 Small, Synthetic). OOD performance declines beyond the generalization ridge.
Layer $\mathbf{I(Z;Y)}$ Test Accuracy (%) All In-Dist. Out-Dist. Layer 1 0.0508 0.00 0.00 0.00 Layer 2 0.0513 0.00 0.00 0.00 Layer 3 0.0619 0.00 0.00 0.00 Layer 4 0.0822 0.00 0.00 0.00 Layer 5 0.0894 0.00 0.00 0.00 Layer 6 0.0989 0.00 0.00 0.00 Layer 7 0.1292 0.70 0.00 2.00 Layer 8 0.2431 18.35 0.00 38.90 Layer 9 0.2810 19.15 1.00 40.30 Layer 10 0.3209 62.45 71.90 53.50 Layer 11 0.3209 72.65 99.90 40.40 Layer 12 0.1402 71.45 100.00 38.60
Information Peak (intermediate layers). $I(Z_{\ell};Y)$ continues to rise through the mid-to-upper layer, typically peaking before the final few blocks. For the GPT-2 Small on Synthetic Arithmetic dataset, the peak reaches $I\!\approx\!0.32$ in layers 10-11, coinciding with a jump from $\approx 0\%$ to $72\%$ ID accuracy and $53\%$ OOD accuracy. These layers appear to play a critical role in synthesizing abstract features that are essential for generalization.
Representational Compression (final layers). Beyond the peak, $I(Z_{\ell};Y)$ decreases, even as in-distribution accuracy approaches $100\%$. The simultaneous drop in OOD accuracy indicates that the final layers tend to memorize training patterns, sacrificing generalization ability.
Notably, this three-phase progression emerges consistently across both GPT-2 Small and GPT-2 Medium, indicating that the observed information dynamics are robust to architectural scale within the same model family.
Generalization Ridge: Memorization-Generalization Trade-off.
The pronounced “information funnel” around intermediate layers reflects a key trade-off between generalization and memorization, which we term the “generalization ridge”. These layers maximize task-relevant information for generalization, while deeper layers increasingly compress and specialize representations, enhancing in-distribution memorization but reducing robustness. This positions intermediate layers as critical control points for managing this trade-off.
Discussion on Model Capacity Threshold
(a) 8 layers: Monotonic Increase
(b) 9 layers: Shallow Peak
(c) Full model:Generalization Ridge
Figure 3: Truncating GPT–2 to 8 layers removes the MI peak; 9–layer variants begin to exhibit a shallow peak, and the full 12–layer model shows a pronounced decline.
We empirically vary the depth of GPT-2 Small models fine-tuned on the CLUTRR dataset to examine how model capacity shapes information dynamics, shown in Figure 3. When the model is truncated to 8 layers, $I(Z;Y)$ increases monotonically—indicating insufficient capacity to develop an information peak. Upon increasing the depth to 9 layers, a shallow peak emerges, signifying the threshold at which the model begins to distinguish generalized features from memorized signals. The full 12-layer model exhibits a clear peak followed by a decline, confirming that generalization ridge emerges only beyond a certain capacity threshold. These results underscore the role of architectural depth in information dynamics and sufficient capacity is required for generalization ridge to emerge.
Discussion on Overfitting Scenario
Figure 4: $I(Z_{\ell};Y)$ rises in final layers during overfitting (LLaMA, ECQA).
To probe overfitting dynamics, we intentionally fine-tuned the model beyond the optimal point. In Figure 4, we observe that $I(Z_{\ell};Y)$ rises again in the final layers—departing from the typical compression phase. This reversal may suggest that the model begins to memorize shortcuts or redundant label-specific noise rather than learning useful information.
5.2 Incremental Information Gain: Information Concentrates at Intermediate Layers
Figure 5: Middle transformer blocks are key to encoding generalizable task-relevant information (GPT-2 Small, Synthetic).
To understand how information accumulates across the network, we compute Incremental Information Gain ($I(\Delta Z_{\ell};Y)$)—the mutual information between each residual transition and the target label embedding. Figure 5 shows the layer-wise gains, revealing that intermediate layers yield the highest information increases. This concentration of information gain highlights their central role in encoding task-relevant features essential for generalization. A case study decoding $\Delta z$ via the LM head is provided in Appendix F, highlighting token-level shifts across layers. The Incremental Information Gain analysis reveals a clear pattern: middle transformer blocks are key to encoding generalizable task-relevant information, thereby forming a generalization ridge. Conversely, later layers contribute little additional predictive signal, and in some cases, actively reduce alignment with the target embeddings. This diminishing contribution in later layers may suggest a shift from general reasoning to memorization. This trend underscores a fundamental trade-off in transformer training dynamics.
6 Residual Scaling Dynamics via Learnable $\beta$ Coefficients
To deepen our understanding of the Generalization Ridge hypothesis, we examine how modulating the contribution of individual transformer blocks affects information flow, and propose a corollary that links layer-wise contribution to generalization performance under distribution shift.
Corollary (Residual Scaling).
Post-peak residual blocks (i.e., layers beyond the generalization ridge) may encode memorized signals. Suppressing these layers’ contribution via learned residual scaling improves out-of-distribution (OOD) generalization, while amplifying them degrades it.
We introduce a residual scaling mechanism with learnable scalar coefficient parameters (Algorithm 1), inspired by prior work on adaptive residual modulation liu2019self ; menghani2024laurel . Transformer architectures inherently employ residual connections to iteratively refine representations. To isolate and quantify the contribution of each transformer blocks, we scale these residual connections with layer-specific scaling factors $\beta_{\ell}\in\mathbb{R}_{\geq 0}$, with definition below. Each $\beta_{\ell}$ controls the strength of the residual contribution from layer $\ell$, enabling the model to adaptively emphasize or suppress specific blocks:
$z^{(\ell)}=z^{(\ell-1)}+\beta_{\ell}\cdot\text{block}^{(\ell)}(z^{(\ell-1)}),\qquad\beta_{\ell}\in\mathbb{R}_{\geq 0}.$
Definition (Residual Scaling Coefficient).
$\beta_{\ell}$ is a learnable scalar parameter associated with transformer layer $\ell$, which modulates the contribution of that layer’s residual output to the model’s forward pass.
Algorithm 1 Residual Scaling: Probing the Contribution of Transformer Blocks to the Generalization–Memorization Trade-off.
1:Pretrained Transformer with $L$ layers, dataset $\mathcal{D}$, learning rate $\eta$
2:Initialize $\beta_{1},\dots,\beta_{L}\leftarrow 1.0$ (trainable); freeze other weights
3:for each training step do
4:     Sample batch $(x,y)\sim\mathcal{D}$
5:     $z^{(0)}\leftarrow\text{Embedding}(x)$
6:     for $\ell=1$ to $L$ do
7:         $r^{(\ell)}\leftarrow\text{block}^{(\ell)}(z^{(\ell-1)})$
8:         $z^{(\ell)}\leftarrow z^{(\ell-1)}+\beta_{\ell}\cdot r^{(\ell)}$
9:     end for
10:     Compute loss $\mathcal{L}(z^{(L)},y)$
11:     Update $\beta_{1},\dots,\beta_{L}$ using gradient descent
12:end for
We freeze model weights and optimize only the residual scaling coefficient parameters $\beta_{\ell}$, which are initialized to 1. These scalars are trained on the (a) in-distribution (ID) split and (b) out-of-distribution (OOD) split. Since no other parameters are updated, the learned $\beta_{\ell}$ values serve as a direct diagnostic of the extent to which each layer should be amplified or attenuated to suit the data regime.
As shown in Table 2, optimizing the residual scaling coefficient parameters on in-distribution results in $\beta_{12}>1$, indicating that in-distribution performance benefits from amplifying the contribution of the final transformer layers—suggestive of their role in memorizing features specific to the training distribution. In contrast, when trained on out-of-distribution (OOD) data, the final layer’s contribution is downweighted ($\beta_{12}<1$), implying that suppressing memorization-heavy components improves generalization. This pattern holds consistently across architectures such as Qwen and LLaMA (Figure 6) and corresponds to improved OOD performance, as shown in Table 3. Together, these findings provide empirical support for the functional specialization of transformer layers: intermediate layers tend to encode generalizable representations, whereas deeper layers increasingly capture memorized signals.
Table 2: $I(Z;Y)$ and residual scaling coefficients $\beta_{\ell}$ for GPT-2 Small. Optimizing on in-distribution data yields $\beta_{12}>1$, while OOD training suppresses the final layer’s contribution with $\beta_{12}<1$.
Layer 1 Layer 2 Layer 3 Layer 4 Layer 5 Layer 6 Layer 7 Layer 8 Layer 9 Layer 10 Layer 11 Layer 12
$I(Z;Y)$ (CLUTRR) 0.0526 0.0474 0.0407 0.0372 0.0319 0.0337 0.0837 0.2688 0.3430 0.3901 0.4514 0.0851
$\beta_{\ell}$ (CLUTRR ID) 0.9998 0.9999 1.0002 1.0002 1.0003 1.0004 0.9993 0.9997 0.9998 0.9993 0.9999 1.0003
$\beta_{\ell}$ (CLUTRR OOD) 1.0005 1.0005 1.0000 1.0000 0.9999 0.9997 1.0002 1.0003 1.0002 0.9995 0.9996 0.9995
$I(Z;Y)$ (Synthetic) 0.0508 0.0513 0.0619 0.0822 0.0894 0.0989 0.1292 0.2431 0.2810 0.3209 0.3209 0.1402
$\beta_{\ell}$ (Synthetic ID) 0.9993 1.0007 0.9995 1.0005 1.0000 0.9999 1.0000 0.9999 0.9995 1.0018 1.0003 1.0001
$\beta_{\ell}$ (Synthetic OOD) 1.0015 1.0003 0.9974 1.0006 1.0010 0.9980 0.9987 0.9974 0.9985 0.9982 0.9981 0.9956
Table 3: Loss and accuracy (%) before and after residual scaling on out-of-distribution (OOD) split. Performance improves following the optimization of residual scaling coefficient parameters.
Dataset Clutrr Synthetic
[…]
GPT-2 Small 6.9655 21.24 6.9636 21.24 2.6533 70.53 2.6485 70.54
Qwen-2.5-0.5B 4.2494 31.01 4.2465 31.01 1.8132 74.31 1.7217 74.55
LLaMA-3.1-8B 1.3209 57.23 1.3193 57.13 1.3927 87.39 1.3155 87.83
Figure 6: Residual scaling coefficients $\beta_{\ell}$ for Qwen-2.5-0.5B and LLaMA-3.1-8B. A similar pattern consistently holds across architectures, with $\beta_{\text{final layer}}$ decreasing under OOD training.
7 Case Study: Attention Dynamics Across Layers
To verify whether generalization ridge layer indeed correspond to semantically meaningful processing, we visualize the attention patterns as a more interpretable signal of where the model focuses. As shown in Figure 7, we visualize attention maps across layers.
Figure 7: Attention map across layers (GPT-2 Small, Synthetic). At the generalization ridge layers, attention becomes more targeted, focusing on predictive tokens.
In early layers (e.g., Layer 1), attention is diffuse and biased toward final tokens—reflecting reliance on position rather than true predict signals.
By generalization ridge layers (e.g., Layer 11), attention becomes more targeted, concentrating on predictive tokens. This shift marks a transition from positional attention to semantically meaningful focus, suggesting that intermediate layers are increasingly capable of isolating task-relevant information from irrelevant content.
In the final layers, attention regresses toward final tokens. This reversion is aligned with the observed decline in $I(Z_{\ell};Y)$. The resurgence of attention towards terminal tokens indicates a potential memorization scenario, where the model re-engages superficial positional strategies, possibly memorizing noise rather than further refining the generalized extraction of predictive signals.
This attention trajectory further supports the generalization ridge hypothesis, highlighting a trade-off between generalization and memorization in the model’s representational strategy.
8 Conclusion
We present InfoRidge, an information-theoretic framework, to analyze how information evolves across layers in transformer language models. By estimating both predictive information and incremental information gain, we systematically characterize the dynamics of task-relevant signal propagation throughout the network. Our findings reveal a consistent generalization ridge emerging in intermediate layers, where mutual information between the hidden representation and the target label reaches its peak. This phenomenon reflects a fundamental trade-off between generalization and memorization as information flows deeper into the model. Residual scaling experiments further corroborate this interpretation, demonstrating the functional specialization of different layers—where intermediate blocks play a key role in supporting generalization, while deeper layers increasingly focus on memorization. Attention flow provides additional insight into where the model concentrates information. Taken together, these results offer a unified perspective on information flow in transformers for natural language generation and point to opportunities for enhancing model architecture, improving interpretability, and increasing robustness to distributional shifts.
9 Limitation
Our analysis focuses on a select set of datasets and model architectures, which may limit the breadth of empirical coverage. The mutual information estimates rely on kernel-based approximations that may be sensitive to representation geometry. Additionally, the residual scaling approach assumes layer-wise independence and does not account for potential inter-layer dependencies during adaptation. Future work could address these limitations by extending the analysis to broader model families and tasks and exploring joint optimization strategies.
Appendix A Mathematical Details for Matrix-Based Information Estimation and Theoretical Foundations
We employ the matrix-based Rényi entropy  [11] to estimate mutual information between model representations and target labels. This method captured sample similarity structure via kernel Gram matrices.
A.1 Matrix-Based Entropy Estimation
Let $U=\{u_{i}\}_{i=1}^{N}\subset\mathbb{R}^{d}$ denote $\ell_{2}$-normalized representations obtained from a specific transformer layer, a residual update, or the embedding of the target label. A Gaussian kernel Gram matrix $G_{U}\in\mathbb{R}^{N\times N}$ is constructed as:
$(G_{U})_{ij}=\exp\left(-\frac{\|u_{i}-u_{j}\|^{2}}{2\sigma^{2}}\right),$
with bandwidth $\sigma=1$. The matrix is then trace-normalized to ensure $\text{tr}(G_{U})=1$.
The matrix-based Rényi entropy of order $\alpha=1$ is defined as:
$H(U)=-\text{tr}(G_{U}\log G_{U}).$
This expression can be interpreted in terms of the eigenvalue spectrum $\{\lambda_{k}\}$ of $G_{U}$, since $G_{U}$ is positive semi-definite and trace-normalized:
$H(U)=-\sum_{k=1}^{N}\lambda_{k}\log\lambda_{k}.$
The entropy thus reflects the dispersion of the eigenvalues. A more uniform spectrum (i.e., higher entropy) suggests more diversity in the representation space, while a sharply peaked spectrum (i.e., low entropy) indicates redundancy or compression.
A.2 Mutual Information Estimation
To estimate the mutual information between two random variables $U$ and $V$, we compute their Gram matrices $G_{U}$ and $G_{V}$, and form the joint similarity matrix via element-wise (Hadamard) product:
$G_{UV}=G_{U}\circ G_{V}.$
After trace-normalization, mutual information is estimated by:
$I(U;V)=H(U)+H(V)-H(U,V),$
where $H(U,V)=-\text{tr}(G_{UV}\log G_{UV})$. The eigenvalue spectrum of $G_{UV}$ governs the joint entropy term; its shape reflects how much of the structure in $U$ and $V$ aligns. A more concentrated spectrum in $G_{UV}$ relative to $G_{U}$ and $G_{V}$ implies stronger dependence and thus higher mutual information.
A.3 Min Wasserstein Generalization Bound [20]
We restate the generalization bound proposed by He et al. [20], which characterizes generalization in terms of the Wasserstein distance between internal representations.
Suppose that the loss function ~$\tilde{\ell}:\mathcal{Y}\times\mathcal{Y}\to\mathbb{R}_{\geq 0}$ is $\rho_{0}$-Lipschitz, and the activation function $\phi_{\ell}:\mathbb{R}\to\mathbb{R}$ is $\rho_{\ell}$-Lipschitz for each $\ell=1,\ldots,L$. Then:
$\small\mathrm{gen}(P_{W|\mathcal{D}_{n}}\!,\!P_{X,Y})\!\leq\!\min_{\ell}\frac{\rho_{0}}{n}\sum_{i=1}^{n}\!\mathbb{E}_{W}\!\left[\!\left(1\!\vee\!\prod_{j=\ell+1}^{L}\!\rho_{j}\|W_{j}\|_{\mathrm{op}}\!\right)\!\cdot\!\mathcal{W}_{1}\!\left(\!P_{T_{\ell},i,Y_{i}|W}(\cdot|W),\!\ P_{T_{\ell},Y|W}(\cdot|W)\!\right)\!\right]$ (3)
This result shows that generalization can be tightly controlled by the Wasserstein distance [21] between internal representations at a specific layer—referred to as the generalization funnel layer.
Connection to Our Work.
The Min Wasserstein Generalization Bound [20] provides a theoretical foundation for our study by characterizing generalization in terms of distributional alignment at an intermediate layer. It motivates our analysis of information flow by suggesting that information is peaked at a specific layer—the generalization funnel. Our InfoRidge framework builds on this insight by quantifying predictive information across layers, and reveals that a specific intermediate layer exhibits peak mutual information and correlates with better generalization performance.
Appendix B Dataset Overview and Statistics
To evaluate information flow and generalization dynamics across model layers, we conduct experiments on three datasets with varying levels of complexity and structure: CLUTRR, ECQA, and a custom-designed Synthetic Arithmetic dataset. Table B.1 summarizes key dataset statistics.
B.1 Dataset Overview
Table 4: Dataset Statistics
Dataset #Train Train Seq. Len #Val Val Seq. Len #Test Test Seq. Len
CLUTRR 9,074 30 2,020 29 1,146 70
ECQA 7,598 21 1,090 21 2,194 21
Synthetic Arithmetic 10,000 9 2,000 9 2,000 9
B.2 CLUTRR
CLUTRR (Compositional Language Understanding and Text-based Relational Reasoning) [25] is a diagnostic benchmark for evaluating relational inference in language models. Each example contains a story describing family relations, and the task is to infer the missing relationship between two entities. The distribution shift stems from clause lengths that are absent in the training set but present during evaluation. We use the task split “gen_train23_test2to10”, where the model is trained on clause lengths 2 and 3 and evaluated on lengths 2 through 10.
B.3 ECQA
ECQA (Explanations for CommonsenseQA) [26] is a commonsense multiple-choice question-answering dataset, where each question is accompanied by 5 answer options.
B.4 Synthetic Arithmetic Dataset
We construct a synthetic diagnostic dataset to disentangle task-relevant signal learning from spurious noise memorization in a controlled setting. Each sample consists of a sequence of 10 symbolic elements, where the signal component follows an arithmetic progression modulo $K$, and the remainder of each element is independently corrupted with random noise. By varying the modulus $K$, we systematically control task complexity and introduce structured shifts in the data distribution.
Synthetic Arithmetic Dataset Construction.
At each position $t$, the signal value is computed as:
$s_{t}=(s_{0}+t\cdot d)\bmod K,\quad\text{with }s_{0}\in[0,K{-}1],\quad d\in[1,K{-}1].$
Each element in the sequence is represented as a string of the form:
[…]
Here $\mathcal{U}_{\text{int}}$ denotes the uniform distribution. The model is trained to predict the signal value of the final (10th) element, using the preceding elements as input context.
For example, with $K=5$, $s_{0}=1$, and $d=2$, a sample might look like:
S1_N42 S3_N77 S0_N18 S2_N56 S4_N90 S1_N11 S3_N65 S0_N23 S2_N37
Each element encodes both a signal (the number following S) and a noise component (the number following N). The target is 4, corresponding to the signal of the final (10th) item in the sequence.
Appendix C Implementation Details
This appendix outlines implementation details in our experiments.
C.1 Prompt Construction
All tasks are cast into a next-token generation format. The model receives a prompt (context) and must generate the correct continuation token. Below are construction strategies and examples for each dataset:
CLUTRR
Each input example in CLUTRR consists of a short narrative describing a set of family relationships, along with a query involving a pair of entities. We construct prompts by concatenating the narrative and a structured natural language question derived from the query tuple. The model is trained to predict the correct relationship as the next token.
Prompt:
Story: [Alice] is [Bob]’s mother. [Bob] is [Charlie]’s father.
Query: What is the relationship between Alice and Charlie? Answer:
Target: grandmother
ECQA (Explanation-augmented Commonsense QA)
Each ECQA instance consists of a multiple-choice question with five candidate answers. We format the prompt by presenting the question followed by all five options (labeled A–E), and conclude with an explicit answer query. The model is trained to predict the correct answer letter as the next token.
Prompt:
Question: What do people usually do at a birthday party?
Options:
A. Sleep
B. Celebrate
C. Cook
D. Exercise
E. Drive
Answer:
Target: B
Synthetic Arithmetic
Each synthetic sample consists of a sequence of 10 symbolic elements, where each element is formatted as S{signal}_N{noise}. The signal values follow an arithmetic progression modulo $K$, and the noise values are independently sampled from a uniform distribution with a fixed range of 100. During training, the modulus $K$ is set to 13. For evaluation, test sequences are generated using values of $K$ from the range $[5,26]$ excluding 13 to simulate a distribution shift. In the residual $\beta_{\ell}$ analysis, we use $K=13$ for in-distribution (ID) training and $K=17$ for out-of-distribution (OOD) training, allowing for a controlled comparison between generalization and memorization behavior. The model receives the first 9 tokens as input and is trained to predict the signal component of the 10th token.
Prompt:
S1_N42 S3_N88 S5_N20 S7_N10 S9_N65 S11_N43 S0_N99 S2_N38 S4_N77
Target: 6
This controlled format enables manipulation of distributional properties by varying the modulus $K$.
C.2 Information Estimation
To estimate mutual information, we subsample between 50 and 200 test examples, with the exact number determined by the model and task, in order to avoid excessive memory usage from large Gram matrices. Full results and statistical significance are reported in Appendix E.
C.3 Residual Scaling with Learnable $\beta_{\ell}$ Parameters
We introduce a vector of learnable scalar weights $\beta=\{\beta_{1},\dots,\beta_{L}\}$ applied to residual connections in a frozen transformer:
$z^{(\ell)}=z^{(\ell-1)}+\beta_{\ell}\cdot\text{Block}^{(\ell)}(z^{(\ell-1)}).$
  • All transformer weights are frozen; only $\beta_{\ell}$ parameters are trained.
  • Each transformer block’s residual output is modulated by $\beta_{\ell}$ via forward hooks.
  • $\beta$ values are initialized to 1 and updated using gradient descent without altering model architecture.
Training and Evaluation Settings
We perform residual $\beta_{\ell}$ training on top of a model that has already been fine-tuned on the target dataset, keeping all previously learned weights frozen and optimizing only the $\beta_{\ell}$ parameters using HuggingFace’s Trainer API. We sweep over batch sizes from 4 to 32 depending on GPU memory constraints, and tune the learning rate between $5\times 10^{-5}$ and $5\times 10^{-6}$. We train for 3–10 epochs based on model capacity and task complexity. Optimization is performed using the AdamW optimizer with a weight decay of 0.01. Evaluation is conducted by comparing the next predicted token with the next ground-truth target token.
Appendix D Compute and Licensing Details
Computing Resources
All experiments were conducted on an NVIDIA RTX A6000 GPU with 48GB of memory.
Model Licenses
The GPT-2 Small [22] and GPT-2 Medium [22] models are released under the MIT License and are available via Hugging Face Transformers. The Qwen-2.5 0.5B [23] model is provided by Alibaba Group under the Apache 2.0 License. The LLaMA-3.1 8B [24] model is made available by Meta under a non-commercial research license and accessed via Hugging Face.
Dataset Licenses
CLUTRR [25] is released under a CC BY 4.0 license as part of EMNLP 2019. The ECQA [26] dataset is also released under a CC BY 4.0 license by its authors as part of EMNLP 2021. The Synthetic Arithmetic dataset is custom-designed by the authors and does not rely on any external or licensed data sources.
All assets were used in compliance with their respective licenses, and no proprietary or restricted resources were employed in our experiments.
Appendix E Full Quantitative Results with Confidence Estimates
To provide a comprehensive view of model behavior across layers, we report the full experimental results for all models and datasets evaluated in this study. This includes predictive information trends ($I(Z;Y)$), incremental information gain ($I(\Delta Z;Y)$) and residual scaling effects ($\beta_{\ell}$), together with statistical significance reporting.
E.1 Predictive Information
To estimate mutual information between hidden representations and target labels, we compute matrix-based mutual information following the methodology outlined in Appendix A. Due to the high memory cost of computing $N\times N$ Gram matrices, we subsample between 50 and 100 test examples per experiment and report results with statistical significance based on multiple random seeds.
(a) GPT-2 Small on CLUTRR
(b) GPT-2 Small on Synthetic Arithmetic
(c) GPT-2 Medium on CLUTRR
(d) GPT-2 Medium on Synthetic Arithmetic
(e) Qwen-2.5-0.5B on ECQA
(f) Qwen-2.5-0.5B on CLUTRR
(g) Qwen-2.5-0.5B on Synthetic Arithmetic
(h) LLaMA-3.1-8B on ECQA
(i) LLaMA-3.1-8B on CLUTRR
(j) LLaMA-3.1-8B on Synthetic Arithmetic
Figure 8: Predictive information $I(Z;Y)$ across different models and datasets exhibits an information peak, indicating a generalization ridge. In cases where the task is too simple relative to model capacity—such as the synthetic arithmetic task with LLaMA—this trend reflects an overfitting regime. Lighter line colors represent later training epochs. Each curve shows the mean across five random seeds (0, 1, 2, 3, 42), and the shaded region denotes a 2-sigma (~96%) confidence interval.
E.2 Incremental Information Gain
In addition to predictive information $I(Z^{(\ell)};Y)$, we compute the incremental information gain $I(\Delta Z^{(\ell)};Y)$ at each layer. This quantity the information contribution for each transformer blocks.
(a) GPT-2 Small on CLUTRR
(b) GPT-2 Small on Synthetic Arithmetic
(c) GPT-2 Medium on CLUTRR
(d) GPT-2 Medium on Synthetic Arithmetic
(e) Qwen-2.5-0.5B on ECQA
(f) Qwen-2.5-0.5B on CLUTRR
(g) Qwen-2.5-0.5B on Synthetic Arithmetic
(h) LLaMA-3.1-8B on ECQA
(i) LLaMA-3.1-8B on CLUTRR
(j) LLaMA-3.1-8B on Synthetic Arithmetic
Figure 9: Incremental information gain $I(\Delta Z;Y)$ across different models and datasets with ~96% CI error bars. Across all models, we observe that the largest incremental information gain consistently occurs in intermediate layers—further supporting the emergence of a generalization ridge.
E.3 Residual Scaling
We introduce a residual scaling mechanism with learnable scalar coefficient parameters, inspired by prior work on adaptive residual modulation [12, 13]. Similar ideas of modulating internal computation have also been explored in parameter-efficient fine-tuning (PEFT) [27], Representation-Efficient Fine-Tuning (REFT) [28] and interpretability-driven control [29, 30, 31, 32]. We present the complete set of residual scaling results, detailing the learned $\beta_{\ell}$ values across all transformer layers. These values reflect the relative contribution of each layer after optimizing the residual scaling coefficients while keeping all other model parameters frozen. $\beta_{\ell}$ are trained on in-distribution (ID) and out-of-distribution (OOD) data. We observe that in the ID setting, later layers tend to receive higher weights, consistent with memorization behavior. In contrast, OOD training consistently downweights final layers and shifts importance toward the middle of the network, aligning with the generalization ridge identified through InfoRidge.
(a) GPT-2 Small on Synthetic Arithmetic
(b) GPT-2 Small on CLUTRR
(c) Qwen-2.5-0.5B on Synthetic Arithmetic
(d) Qwen-2.5-0.5B on CLUTRR
(e) LLaMA-3.1-8B on Synthetic Arithmetic
(f) LLaMA-3.1-8B on CLUTRR
Figure 10: Residual scaling coefficients $\beta_{\ell}$ across all transformer layers. ID training emphasizes later layers, while OOD training shifts weight toward middle layers, aligning with the generalization ridge observed via InfoRidge. Each curve shows the mean across five random seeds (0, 1, 2, 3, 42), and the shaded region denotes 1-sigma error bar.
Appendix F Incremental Information Gain Case Study
To further illustrate how intermediate layers contribute to generalization, we analyze the semantic content introduced by residual transitions at different depths. Specifically, we decode the residual transition $\delta z$ at each layer using the language modeling (LM) head, projecting the incremental representation back into token space.
This analysis allows us to inspect the linguistic shift introduced by each transformer block in isolation, and to assess whether the changes correspond to task-relevant predictions or superficial noises.
F.1 Methodology
For a fixed input, we compute residual transitions $\delta z^{(\ell)}$ at each layer and pass them through the model’s final linear projection (LM head) followed by a softmax. We record the top predicted tokens and their probabilities, and it reflects the directional change applied by layer $\ell$.
F.2 Observations
We decode residual transitions across layers and report the top shifted tokens by projecting $\Delta Z^{(\ell)}$ through the LM head. These shifts provide insight into how each layer modifies the model’s internal prediction trajectory.
Layer 9 → 10: The top shifted tokens include ĠNine, Ġ10, Ġ4, Ġ8, and 8, which are all numerically aligned with the target prediction space in the synthetic arithmetic task. This indicates that the model is beginning to refine task-relevant numerical features at this depth.
Layer 10 → 11: The shifted tokens become partially diluted, featuring punctuation and less informative symbols such as ,, Ġ, and ., alongside occasional task-relevant entries like Ġ12 and Ġ4. This indicates a transitional phase where the model continues to refine meaningful task-relevant features, yet begins to exhibit noise from frequent but semantically uninformative, such as punctuation, tokens.
Layer 11 → 12: At the final layer transition, the top 5 shifted tokens become largely uninterpretable, including Ġcanaver, soDeliveryDate, enegger, 76561, and ILA. Conversely, the most negatively shifted tokens—Ġ4, Ġ3, Ġ5, Ġ6, Ġ1—correspond to plausible numerical predictions that were actively suppressed. This supports the hypothesis that final layers may overwrite generalizable abstractions with memorized or noise signals.
These patterns are consistent with our mutual information analysis, which identifies intermediate layers as better semantically aligned with the prediction target—forming a ridge of generalization.
F.3 Implication
These decoding results reinforce our interpretation of the generalization ridge: intermediate layers contribute the most semantically informative updates to the model’s representation. The residual transitions thus serve as a useful lens for understanding how and where semantic meaning is introduced during forward propagation.

### v2

@@
 Abstract
-Transformer-based language models have achieved state-of-the-art performance in natural language generation (NLG) tasks, yet their internal mechanisms for synthesizing task-relevant information remain insufficiently understood.
+Transformer-based language models have achieved state-of-the-art performance in natural language generation (NLG), yet their internal mechanisms for synthesizing task-relevant information remain insufficiently understood.
 While prior studies suggest that intermediate layers often yield more generalizable representations than final layers, how this generalization ability emerges and propagates across layers during training remains unclear.
@@
 To further investigate this phenomenon, we introduce residual scaling coefficients—trainable scalar parameters applied to each residual block—which serve as functional probes for assessing the relative importance of individual transformer layers.
-These coefficients reveal that, under distribution shift, models downweight final layers and increasingly rely on ridge layers, highlighting their role in generalization.
+These coefficients reveal that, under distribution shift, models downweight final layers and increasingly rely on intermediate layers, highlighting their role in generalization.
 Together, these findings offer new insights into the internal mechanisms of transformers and underscore the critical role of intermediate layers in supporting generalization.
 1 Introduction
-Transformer-based language models have achieved remarkable performance in natural language generation (NLG) tasks such as summarization, machine translation, and dialogue generation vaswani2017attention ; dong2022survey .
+Transformer-based language models generate text by predicting tokens autoregressively, and they have achieved remarkable performance across a wide range of natural-language uses (Vaswani et al., 2017; Dong et al., 2022).
 Nevertheless, we lack a rigorous understanding of how these models acquire and synthesize task-relevant information during training.
-A growing body of research has shown that intermediate layers in deep neural networks often surpass final layers in terms of representational quality and generalization performance liu2019linguistic ; voita2019bottom ; ansuini2019intrinsic ; ahrens2023read ; uselis2025intermediate .
-In language models, intermediate layers often encode richer semantic and more robust features than final layers fan2024not ; jin2024exploring ; skean2025layer .
+A growing body of research has shown that intermediate layers in deep neural networks often surpass final layers in terms of representational quality and generalization performance (Liu et al., 2019b; Voita et al., 2019; Ansuini et al., 2019; Ahrens et al., 2023; Uselis & Oh, 2025).
+In language models, intermediate layers often encode richer semantic and more robust features than final layers (Fan et al., 2024; Jin et al., 2024; Skean et al., 2025).
 However, questions still remain: In NLG, how does information evolve across layers during training, and how are different layers of the network functionally organized to support generalization versus memorization?
 To investigate these questions, we propose InfoRidge, an information-theoretic framework, to analyze information flow in language models.
-Building on matrix-based mutual information estimation giraldo2014measures , our approach quantifies how predictive signals transform across layers during training.
-We center our analysis on two complementary quantities.
-The first is the predictive information at each layer, defined as the mutual information $I(Z_{\ell};Y)$ between the hidden representation $Z_{\ell}$ at layer $\ell$ and the next-token label $Y$, which reflects how much task-relevant information is preserved.
-The second is the incremental information gain, denoted as $I(\Delta Z_{\ell};Y)$, where $\Delta Z_{\ell}$ is the residual changes between successive layer $\ell$ and $\ell-1$.
-This measures the additional predictive information introduced by each transformer block.
-Using these tools, we uncover a non-monotonic trend: predictive information rises through early and middle layers, peaks in the upper-middle, and then declines in later layers.
-We term this phenomenon the generalization ridge, indicating a structural division of labor across depth.
-Intermediate layers appear to specialize in encoding abstract, generalizable features, while later layers increasingly reorient representations toward task-specific alignment and memorization.
-By further analyzing incremental information gain, we find that certain middle layers introduce the largest increases in predictive information, marking them as key contributors to the emergence of the ridge.
-Such layer-wise specialization offers insight into how generalization and memorization are distributed within the network.
-To further validate this interpretation, we introduce residual scaling coefficients—learnable scalar parameters $\beta_{\ell}$ applied to each residual block—while keeping all other model weights fixed, drawing inspiration from prior work on layer-wise adaptation liu2019self ; menghani2024laurel .
+Building on matrix-based mutual information estimation (Giraldo et al., 2014), our approach quantifies how predictive signals transform across layers during training.
+We center our analysis on two complementary quantities: predictive information, defined as the mutual information $I(Z_{\ell};Y)$ between the hidden representation $Z_{\ell}$ at layer $\ell$ and the next-token label $Y$, reflecting how much task-relevant information is preserved; incremental information gain, denoted as $I(\Delta Z_{\ell};Y)$, where $\Delta Z_{\ell}$ is the residual changes between successive layer $\ell$ and $\ell-1$, measuring the additional predictive information introduced by each transformer block.
+Using InfoRidge, we uncover a non-monotonic trend: predictive information rises through the early and middle layers, peaks in the upper-middle layers, and then declines in the later layers.
+We name this phenomenon the generalization ridge, where the model encodes the more generalizable task-relevant information.
+This ridge marks a structural division of labor: intermediate layers concentrate generalizable features that transfer across distributions, while later layers increasingly specialize in task-specific memorization.
+Incremental information gain further shows that ridge layers introduce the largest increases in predictive information, marking them as key contributors to the emergence of the ridge.
+This analysis directly connects the information peak to generalizable behavior and clarifies how generalization and memorization are distributed across depth.
+To further validate this interpretation, we introduce residual scaling coefficients—learnable scalar parameters $\beta_{\ell}$ applied to each residual block—while keeping all other model weights frozen, drawing inspiration from prior work on layer-wise adaptation (Liu et al., 2019a; Menghani et al., 2024).
 A higher $\beta_{\ell}$ value indicates greater reliance on the corresponding layer’s output during prediction.
@@
 Under in-distribution training, deeper layers retain higher residual weights, reflecting the model’s reliance on memorized, task-specific features.
-When evaluated under distribution shift, models reduce reliance on late layers and increase reliance on the ridge, further supporting its role in generalization.
+When evaluated under distribution shift, models have lower reliance on the final layer and shifts importance toward the intermediate layers.
 To understand the formation of the generalization ridge, we analyze both attention patterns and model capacity.
-Attention flow analysis shows that ridge layers attend more directly to semantically informative tokens, suggesting that the ridge is not only a statistical peak but also functionally aligned with meaningful computation.
-In contrast, deeper layers increasingly focus on positional tokens, consistent with a shift toward memorization.
-Beyond functional behavior, we also investigate the conditions under which the ridge emerges.
+Attention analysis shows ridge layers attend to tokens that capture broadly useful features, aligning with the information peak.
+Beyond attention patterns, we also investigate the conditions under which the ridge emerges.
 Our depth ablation results show that the ridge only emerges beyond a certain depth threshold.
@@
 Our work provides a unified perspective on how predictive information is structured across depth in transformer-based language models for natural language generation tasks:
-  1. 1.
-    We propose InfoRidge, an information-theoretic framework, to quantify both predictive information and incremental information gain across layers, offering a dynamic view of how task-relevant signals emerge and evolve during training.
-  2. 2.
-    Through this framework, we identify the generalization ridge—a mid-layer peak in predictive information—revealing a transition between generalization and memorization.
-  3. 3.
-    We propose residual scaling coefficients as functional probes for layer importance, uncovering a shift in reliance from final layers to intermediate layers under distribution shift, consistent with the ridge’s role in supporting generalization.
-These findings suggest that transformer depth reflects functional stratification: intermediate layers encode core abstractions for generalization, while final layers increasingly tailor representations for memorization—often at the cost of robustness.
-This insight informs the design of more interpretable and adaptable architectures under distribution shift.
+1.
+Our work tracks the evolution of predictive information throughout training, establishing a clear connection between predictive information flow and generalization.
+It reveals a non-monotonic peak in the middle layers, which we refer to as the generalization ridge.
+This pattern reflects a meaningful transition in representational focus and aligns with stronger generalization behavior.
+2. We introduce InfoRidge, an information-theoretic framework that applies matrix-based mutual information estimation to autoregressive language models to analyze information flow.
+3.
+We introduce residual scaling coefficients as trainable indicators of how models shift representational focus across layers during training, providing a causal, adaptive measure of model behavior under distribution shifts.
 2 Related Work
-Understanding how information is encoded and transformed across layers in deep models has been widely studied through probing classifiers alain2016understanding , attention flow analysis vig2019analyzing , and information-theoretic tools such as the information bottleneck shwartz2017opening , mutual information estimation goldfeld2019estimating , and matrix-based entropy methods giraldo2014measures .
-These methods provide various lenses to quantify representational capacity, abstraction, and invariance across layers.
-A growing body of research has shown that intermediate layers in deep networks often outperform final layers in terms of representational quality and task performance ansuini2019intrinsic ; yosinski2014transferable ; uselis2025intermediate ; ahrens2023read ; ando2023use .
-In language models, mid-depth layers tend to capture richer semantic or robust features than output layers liu2019linguistic ; voita2019bottom ; jin2024exploring ; fan2024not .
-These findings challenge the assumption that deeper layers always yield better representations, revealing instead a peak in representational utility at intermediate depths.
-This pattern holds across settings such as transfer learning yosinski2014transferable , continual learning ahrens2023read , and out-of-distribution generalization uselis2025intermediate .
-Recent work by Skean et al.
-skean2025layer proposes a unified framework that evaluates representation quality using entropy, curvature, and invariance.
-Collectively, these results highlight the importance of intermediate layers for abstraction and generalization, motivating our analysis of layer-wise information dynamics through an information-theoretic lens.
-Differently, we focus on information flow for next token generation, specifically examining mutual information between the last input token and the target label.
-Our goal is to understand how generalization and memorization dynamics evolve during training from information-theoretic perspective.
+Understanding how information is encoded and transformed across layers has been studied through probing classifiers (Alain & Bengio, 2016), attention flow (Vig & Belinkov, 2019), and information-theoretic approaches such as the information bottleneck (Shwartz-Ziv & Tishby, 2017), mutual information estimation (Goldfeld, 2019), and matrix-based entropy (Giraldo et al., 2014), offering different lenses to quantify representational capacity, abstraction, and invariance across layers.
+A growing body of research has shown that intermediate layers in deep networks often outperform final layers in terms of representational quality and task performance (Ansuini et al., 2019; Yosinski et al., 2014; Uselis & Oh, 2025; Ahrens et al., 2023; Ando et al., 2023).
+In language models, mid-depth layers tend to capture richer semantic or robust features than output layers (Liu et al., 2019b; Voita et al., 2019; Jin et al., 2024; Fan et al., 2024).
+Transformer representations have long been observed to follow a structured progression from syntactic to semantic information, as shown by classical probing studies on linguistic knowledge and the reconstruction of the NLP pipeline (Liu et al., 2019b; Tenney et al., 2019).
+These findings challenge the assumption that deeper layers always yield better representations.
+This pattern holds across settings such as transfer learning (Yosinski et al., 2014), continual learning (Ahrens et al., 2023), and out-of-distribution generalization (Uselis & Oh, 2025).
+Furthermore, recent work has evaluated representation quality using entropy, curvature, and invariance (Skean et al., 2025), while other studies have examined embedding drift and representational geometry (Merchant et al., 2020; Dar et al., 2022), analyzed memorization and factual recall (Haviv et al., 2022; Yu et al., 2023), and introduced causal perspectives on layer importance through mediation analysis and targeted interventions (Vig et al., 2020; Meng et al., 2022).
+Training-dynamics studies investigated how earlier models develop and refine semantic features across depth, providing an additional perspective on the evolution of layer-wise representations (Merchant et al., 2020; Kumar et al., 2023).
+Correlational probes (e.g., linear probes (Alain & Bengio, 2016)) measure only whether a feature can be decoded from a representation, which reflects correlation but not causal influence.
+In contrast, causal methods intervene on internal activations to test how changing a component alters the model’s prediction, thereby identifying true causal contribution rather than mere feature presence.
+Our residual-scaling approach aligns with this causal perspective at the layer level.
+However, the underlying causes and functional role of this phenomenon remain only partially understood, motivating further investigation.
+Our work addresses this gap by tracing the evolution of predictive information throughout training and establishing a clear connection between predictive information flow and generalization.
+We reveal a consistent non-monotonic peak in the middle layers—termed the generalization ridge—that reflects a meaningful transition in representational focus and aligns with stronger generalization behavior.
+Additionally, unlike prior work focused on classification tasks, we extend the analysis of information flow to generation tasks by quantifying the mutual information between hidden states and the next token.
+This enables us to understand how generalization and memorization dynamics evolve for next-token generation setting during training from an information-theoretic perspective.
 3 InfoRidge: Information Estimation Framework
 Figure 1: Overview of InfoRidge.
-(1) Extract internal representations at each layer and compute residual changes between successive layers.
-(2) Extract the target token embedding.
-(3) Estimate predictive information and incremental information gain.
+(1) Step 1: Extract internal representations at each layer and compute residual changes between successive layers.
+(2) Step 2: Extract the target token embedding.
+(3) Step 3: Using these representations, we estimate the layer-wise predictive information $I(Z_{\ell};Y)$ and the incremental information gain $I(\Delta Z_{\ell};Y)$, which jointly characterize how information evolves across depth.
 We propose InfoRidge, an information-theoretic framework that uses mutual information to quantify how predictive information propagates through transformers layers during training in NLG.
 Motivating Insight.
-Prior work has shown that internal representations in deep neural networks tend to align most closely with the true data distribution at an intermediate layer he2024information .
-By employing the Wasserstein distance villani2008optimal , this alignment is shown to reach a minimum at a specific depth—referred to as the generalization funnel layer.
-At this point, the Min Wasserstein Generalization Bound he2024information ensures that the upper bound on the generalization gap—defined as the expected difference between the population and empirical risks—is minimized.
+Prior work has shown that internal representations in deep neural networks tend to align most closely with the true data distribution at an intermediate layer (He et al., 2024).
+By employing the Wasserstein distance (Villani et al., 2008), this alignment is shown to reach a minimum at a specific depth—referred to as the generalization funnel layer.
+At this point, the Min Wasserstein Generalization Bound (He et al., 2024) ensures that the upper bound on the generalization gap—defined as the expected difference between the population and empirical risks—is minimized.
 This highlights the critical role of intermediate layers in supporting generalization.
+Research Question.
+Despite the insight from prior studies of deep neural networks in classification, it remains unclear whether this generalizes to Transformer-based language models in natural language generation (NLG).
+This motivates our central question: In NLG, how does information evolve across layers during training, and how are different layers of the network functionally organized to support generalization versus memorization?
 Hypothesis.
-Building on this insight, we hypothesize that there exists a specific intermediate transformer layer that encodes the most generalizable representations for next-token prediction, characterized by maximal mutual information with the target label.
+Building on the insight, we hypothesize that there exists a specific intermediate transformer layer that encodes the most generalizable representations for next-token prediction, characterized by maximal mutual information with the target label.
 Hypothesis: Generalization Ridge.  There exists an intermediate layer $\ell^{*}\in\{1,\dots,L\}$ such that the mutual information between the hidden state and the target label peaks at that layer:
@@
 Matrix-Based Mutual Information.
-We apply the matrix-based framework giraldo2014measures to estimate mutual information.
+We apply the matrix-based framework (Giraldo et al., 2014) to estimate mutual information.
 Let $\mathcal{U}$ be a random variable, from which we draw a set of vectors $U=\{\mathbf{u}_{i}\}_{i=1}^{N}\subset\mathbb{R}^{d}$.
@@
 Models.
-We evaluate four models: GPT–2 Small (117M) radford2019language , GPT–2 Medium (345M) radford2019language , Qwen–2.5 0.5B qwen2 , and LLaMA 3.1 8B meta_llama_3.1_8b .
+We evaluate four models: GPT–2 Small (117M) (Radford et al., 2019), GPT–2 Medium (345M) (Radford et al., 2019), Qwen–2.5 0.5B (Yang et al., 2024), and LLaMA 3.1 8B (Meta AI, 2024).
 All models are fine-tuned on NLG tasks.
@@
 Datasets.
-We assess model behavior across three tasks reformulated as NLG problems: CLUTRR (sinha2019clutrr, ), a relational reasoning benchmark; ECQA (aggarwal2021explanations, ), a commonsense QA benchmark; and Synthetic Arithmetic, a controlled dataset designed to disentangle task-relevant signal from noise.
-Dataset statistics and implementation details are provided in Appendix B and C.
-[…]
+We assess model behavior across three tasks casted into NLG problems: CLUTRR (Sinha et al., 2019), a relational reasoning benchmark; ECQA (Aggarwal et al., 2021), a commonsense QA benchmark; and Synthetic Arithmetic, a controlled dataset designed to disentangle task-relevant signal from noise.
+Dataset and implementation details are in Appendix B and C.
+Synthetic Arithmetic Dataset Construction.
+We construct a synthetic dataset to separate signal learning from noise memorization.
+Each input is a sequence of 10 elements, where the signal follows an arithmetic progression modulo $K$, computed as $s_{t}=(s_{0}+t\cdot d)\bmod K$ with $s_{0}\in[0,K{-}1]$ and $d\in[1,K{-}1]$.
+Each element takes the form S{signal}_N{noise}, where noise is sampled from $\mathcal{U}_{\text{int}}(0,\texttt{noise\_range-1})$ ($\mathcal{U}_{\text{int}}$ denotes the uniform distribution).
+The model is trained to predict the signal value of the final (10th) element using the preceding elements as input context.
+For example, with $K=5$, $s_{0}=1$, and $d=2$, a sample input might be S1_N42 S3_N77...S2_N37, with the target signal being 4.
+By varying $K$, we can induce structured distribution shifts.
 5 Generalization Ridge: Layer-wise Mutual Information Training Dynamics
@@
 5.1 Predictive Information: Information Peaks at Intermediate Layers
-We investigate how predictive information—defined as the mutual information between hidden representations and target labels—evolves across transformer layers.
-Specifically, we calculate the matrix–based mutual information $I(Z_{\ell};Y)$ at each transformer layer $\ell$, capturing the amount of label-relevant information encoded during forward propagation.
+We investigate how predictive information—defined as the mutual information between hidden representations and target labels—evolves across the depth of transformer models.
+Specifically, for each layer $\ell$, we compute the matrix-based mutual information $I(Z_{\ell};Y)$ between the hidden state $Z_{\ell}$ and the next-token ground truth $Y$.
+This quantity measures how much task-relevant signal is retained in the representation as it propagates through the network.
+By tracing $I(Z_{\ell};Y)$ across layers, we obtain a layer-wise trajectory of information flow, which reveals not only where predictive content is preserved but also how it is transformed or diminished as the model processes.
 (a) GPT-2 Small on Synthetic
@@
 [^1] The in distribution split is generated with $K_{\text{id}}{=}13$; out of distribution splits use a uniformly–sampled $K_{\text{ood}}\in\{5,\dots,25\}\setminus\{13\}$.
-Additional results are presented in Appendix E.
+Additional results are in Appendix E.
 Three-phase information dynamics.
 Across models and tasks, predictive information curves exhibit a consistent three-phase pattern:
+Table 1: Information Dynamics and Layer-wise Performance (GPT-2-S, Synthetic). OOD performance declines beyond the generalization ridge.
+Layer $\mathbf{I(Z;Y)}$ Test Accuracy (%) All In-Dist.
+Out-Dist.
+Layer 1 0.0782 0.00 0.00 0.00 Layer 2 0.0776 0.00 0.00 0.00 Layer 3 0.0907 0.00 0.00 0.00 Layer 4 0.1156 0.00 0.00 0.00 Layer 5 0.1222 0.00 0.00 0.00 Layer 6 0.1275 0.00 0.00 0.00 Layer 7 0.1552 0.75 0.00 1.50 Layer 8 0.2594 19.45 0.00 40.70 Layer 9 0.2937 20.60 0.60 42.00 Layer 10 0.3295 64.25 72.90 56.90 Layer 11 0.3296 71.65 100.00 42.90 Layer 12 0.1501 70.45 100.00 41.40
 Progressive Accrual (early layers).
@@
 This aligns with the near-zero accuracy observed in Table 1 for these layers.
-Table 1: Information Dynamics and Layer-wise Performance (GPT-2 Small, Synthetic). OOD performance declines beyond the generalization ridge.
-Layer $\mathbf{I(Z;Y)}$ Test Accuracy (%) All In-Dist.
-Out-Dist.
-Layer 1 0.0508 0.00 0.00 0.00 Layer 2 0.0513 0.00 0.00 0.00 Layer 3 0.0619 0.00 0.00 0.00 Layer 4 0.0822 0.00 0.00 0.00 Layer 5 0.0894 0.00 0.00 0.00 Layer 6 0.0989 0.00 0.00 0.00 Layer 7 0.1292 0.70 0.00 2.00 Layer 8 0.2431 18.35 0.00 38.90 Layer 9 0.2810 19.15 1.00 40.30 Layer 10 0.3209 62.45 71.90 53.50 Layer 11 0.3209 72.65 99.90 40.40 Layer 12 0.1402 71.45 100.00 38.60
 Information Peak (intermediate layers).
 $I(Z_{\ell};Y)$ continues to rise through the mid-to-upper layer, typically peaking before the final few blocks.
-For the GPT-2 Small on Synthetic Arithmetic dataset, the peak reaches $I\!\approx\!0.32$ in layers 10-11, coinciding with a jump from $\approx 0\%$ to $72\%$ ID accuracy and $53\%$ OOD accuracy.
+For the GPT-2 Small on Synthetic Arithmetic dataset, the peak reaches $I\!\approx\!0.33$ in layers 10-11, coinciding with a jump from $\approx 0\%$ to $73\%$ ID accuracy and $57\%$ OOD accuracy.
 These layers appear to play a critical role in synthesizing abstract features that are essential for generalization.
@@
 This positions intermediate layers as critical control points for managing this trade-off.
-Discussion on Model Capacity Threshold
+Generalization Ridge Emerges Beyond a Depth Threshold
 (a) 8 layers: Monotonic Increase
 (b) 9 layers: Shallow Peak
-(c) Full model:Generalization Ridge
+(c) Full model: Gen Ridge
 Figure 3: Truncating GPT–2 to 8 layers removes the MI peak; 9–layer variants begin to exhibit a shallow peak, and the full 12–layer model shows a pronounced decline.
@@
 Upon increasing the depth to 9 layers, a shallow peak emerges, signifying the threshold at which the model begins to distinguish generalized features from memorized signals.
-The full 12-layer model exhibits a clear peak followed by a decline, confirming that generalization ridge emerges only beyond a certain capacity threshold.
-These results underscore the role of architectural depth in information dynamics and sufficient capacity is required for generalization ridge to emerge.
-Discussion on Overfitting Scenario
+The full 12-layer model exhibits a clear peak followed by a decline, confirming that the generalization ridge emerges only beyond a certain capacity threshold.
+These results underscore the role of architectural depth in information dynamics and sufficient capacity is required for the generalization ridge to emerge.
+Overfitting Adds Memorization in Final Layers
 Figure 4: $I(Z_{\ell};Y)$ rises in final layers during overfitting (LLaMA, ECQA).
 To probe overfitting dynamics, we intentionally fine-tuned the model beyond the optimal point.
-In Figure 4, we observe that $I(Z_{\ell};Y)$ rises again in the final layers—departing from the typical compression phase.
-This reversal may suggest that the model begins to memorize shortcuts or redundant label-specific noise rather than learning useful information.
+In Figure 4, we observe that the Predictive Information $I(Z_{\ell};Y)$ rises again in the final layers—departing from the typical compression phase expected at this stage.
+Before overfitting, the top layers largely behave as pretrained decoders, focusing on surface-level patterns such as token co-occurrence, syntactic templates, or corpus biases, which are not strongly aligned with the task label—hence their initially low $I(Z_{\ell};Y)$.
+After overfitting, however, the later-layer rise reflects a shift toward memorization, where the model begins to encode superficial shortcuts or redundant label-specific noise rather than useful task information.
+Semantically Important Attention Peaks Where Information Peaks
+This discussion aims to provide an intuitive illustration of our generalization ridge hypothesis, highlighting how attention to task-relevant signal tokens shifts across layers (the generalizable information).
+Specifically, we computed average signal attention across layers, identified the layer with peak signal attention, and compared it to final-layer signal and last-token attention, results are in Table 2.
+Signal tokens are defined as task-relevant tokens that the model must focus on to solve the task—for the synthetic dataset, these are tokens that appear after the character ‘S’; for CLUTRR, kinship-related entities; and for ECQA, the token corresponding to the correct answer option.
+This allows us to quantify where in the network attention to semantically important information is concentrated.
+Our findings show a pattern that supports our hypothesis: (1) Mid-to-late layers peak in signal attention, coinciding with the predictive information ridge (Figure 2), indicating where generalizable representations are strongest.
+(2) Final layers show reduced attention to signal tokens, suggesting a shift toward memorization to specific data point rather than predictive information abstraction.
+For example, in Qwen-2.5-0.5B on ECQA, signal attention peaks at Layer 17 (0.2458) but drops to 0.0066 in the final layer, where last-token attention dominates (0.1697).
+Additional results are in Appendix F.
+Table 2: Average attention statistics: (1) average attention scores over all tokens, (2) average attention to signal tokens, (3) the maximum signal attention and its corresponding layer, (4) signal token attention in the final layer and (5) last token attention in the final layer.
+Model Dataset Avg. Attn (All) Avg. Attn (Signal) Layer w/ Highest Avg. Signal Final Avg. Signal Final Avg. Last
+GPT-2 Small Synthetic 0.0227 0.0483 10 (0.0758) 0.0573 0.0400
+GPT-2 Medium CLUTRR 0.0086 0.0155 19 (0.0453) 0.0147 0.0889
+Qwen-2.5-0.5B ECQA 0.0225 0.0257 17 (0.2458) 0.0066 0.1697
+LLaMA-3.1-8B ECQA 0.0220 0.0379 21 (0.1276) 0.0168 0.3904
+*We remove the first token attention score to mitigate attention sink effects.
 5.2 Incremental Information Gain: Information Concentrates at Intermediate Layers
-Figure 5: Middle transformer blocks are key to encoding generalizable task-relevant information (GPT-2 Small, Synthetic).
 To understand how information accumulates across the network, we compute Incremental Information Gain ($I(\Delta Z_{\ell};Y)$)—the mutual information between each residual transition and the target label embedding.
-Figure 5 shows the layer-wise gains, revealing that intermediate layers yield the highest information increases.
-This concentration of information gain highlights their central role in encoding task-relevant features essential for generalization.
-A case study decoding $\Delta z$ via the LM head is provided in Appendix F, highlighting token-level shifts across layers.
+As shown in Figure 5, the resulting layer-wise gains reveal that intermediate layers yield the highest information increases.
+This concentration of information gain further underscores their central role in encoding those task-relevant features that are essential for supporting generalization.
+For additional illustration, a detailed case study that decodes $\Delta z$ via the LM head is provided in Appendix G, highlighting fine-grained token-level shifts and variations observed across layers.
+Figure 5: Middle blocks are key to encoding generalizable information (GPT-2-S, Synthetic).
 The Incremental Information Gain analysis reveals a clear pattern: middle transformer blocks are key to encoding generalizable task-relevant information, thereby forming a generalization ridge.
 Conversely, later layers contribute little additional predictive signal, and in some cases, actively reduce alignment with the target embeddings.
-This diminishing contribution in later layers may suggest a shift from general reasoning to memorization.
+This diminishing contribution in later layers may suggest a shift from general reasoning to memorization of training-specific patterns.
 This trend underscores a fundamental trade-off in transformer training dynamics.
@@
 Post-peak residual blocks (i.e., layers beyond the generalization ridge) may encode memorized signals.
-Suppressing these layers’ contribution via learned residual scaling improves out-of-distribution (OOD) generalization, while amplifying them degrades it.
-We introduce a residual scaling mechanism with learnable scalar coefficient parameters (Algorithm 1), inspired by prior work on adaptive residual modulation liu2019self ; menghani2024laurel .
+Suppressing these layers’ contribution via learned residual scaling improves out-of-distribution generalization, while amplifying them degrades it.
+We introduce a residual scaling mechanism with learnable scalar coefficient parameters (Algorithm 1), inspired by prior work on adaptive residual modulation (Liu et al., 2019a; Menghani et al., 2024).
 Transformer architectures inherently employ residual connections to iteratively refine representations.
-To isolate and quantify the contribution of each transformer blocks, we scale these residual connections with layer-specific scaling factors $\beta_{\ell}\in\mathbb{R}_{\geq 0}$, with definition below.
+To isolate and quantify the contribution of each transformer block, we scale these residual connections with layer-specific scaling factors $\beta_{\ell}\in\mathbb{R}_{\geq 0}$, with definition below.
 Each $\beta_{\ell}$ controls the strength of the residual contribution from layer $\ell$, enabling the model to adaptively emphasize or suppress specific blocks:
@@
 We freeze model weights and optimize only the residual scaling coefficient parameters $\beta_{\ell}$, which are initialized to 1.
-These scalars are trained on the (a) in-distribution (ID) split and (b) out-of-distribution (OOD) split.
-Since no other parameters are updated, the learned $\beta_{\ell}$ values serve as a direct diagnostic of the extent to which each layer should be amplified or attenuated to suit the data regime.
-As shown in Table 2, optimizing the residual scaling coefficient parameters on in-distribution results in $\beta_{12}>1$, indicating that in-distribution performance benefits from amplifying the contribution of the final transformer layers—suggestive of their role in memorizing features specific to the training distribution.
-In contrast, when trained on out-of-distribution (OOD) data, the final layer’s contribution is downweighted ($\beta_{12}<1$), implying that suppressing memorization-heavy components improves generalization.
-This pattern holds consistently across architectures such as Qwen and LLaMA (Figure 6) and corresponds to improved OOD performance, as shown in Table 3.
-Together, these findings provide empirical support for the functional specialization of transformer layers: intermediate layers tend to encode generalizable representations, whereas deeper layers increasingly capture memorized signals.
-Table 2: $I(Z;Y)$ and residual scaling coefficients $\beta_{\ell}$ for GPT-2 Small.
-Optimizing on in-distribution data yields $\beta_{12}>1$, while OOD training suppresses the final layer’s contribution with $\beta_{12}<1$.
-Layer 1 Layer 2 Layer 3 Layer 4 Layer 5 Layer 6 Layer 7 Layer 8 Layer 9 Layer 10 Layer 11 Layer 12
-$I(Z;Y)$ (CLUTRR) 0.0526 0.0474 0.0407 0.0372 0.0319 0.0337 0.0837 0.2688 0.3430 0.3901 0.4514 0.0851
-$\beta_{\ell}$ (CLUTRR ID) 0.9998 0.9999 1.0002 1.0002 1.0003 1.0004 0.9993 0.9997 0.9998 0.9993 0.9999 1.0003
-$\beta_{\ell}$ (CLUTRR OOD) 1.0005 1.0005 1.0000 1.0000 0.9999 0.9997 1.0002 1.0003 1.0002 0.9995 0.9996 0.9995
-$I(Z;Y)$ (Synthetic) 0.0508 0.0513 0.0619 0.0822 0.0894 0.0989 0.1292 0.2431 0.2810 0.3209 0.3209 0.1402
-$\beta_{\ell}$ (Synthetic ID) 0.9993 1.0007 0.9995 1.0005 1.0000 0.9999 1.0000 0.9999 0.9995 1.0018 1.0003 1.0001
-$\beta_{\ell}$ (Synthetic OOD) 1.0015 1.0003 0.9974 1.0006 1.0010 0.9980 0.9987 0.9974 0.9985 0.9982 0.9981 0.9956
-Table 3: Loss and accuracy (%) before and after residual scaling on out-of-distribution (OOD) split. Performance improves following the optimization of residual scaling coefficient parameters.
-Dataset Clutrr Synthetic
-[…]
-GPT-2 Small 6.9655 21.24 6.9636 21.24 2.6533 70.53 2.6485 70.54
-Qwen-2.5-0.5B 4.2494 31.01 4.2465 31.01 1.8132 74.31 1.7217 74.55
-LLaMA-3.1-8B 1.3209 57.23 1.3193 57.13 1.3927 87.39 1.3155 87.83
-Figure 6: Residual scaling coefficients $\beta_{\ell}$ for Qwen-2.5-0.5B and LLaMA-3.1-8B.
-A similar pattern consistently holds across architectures, with $\beta_{\text{final layer}}$ decreasing under OOD training.
-7 Case Study: Attention Dynamics Across Layers
-To verify whether generalization ridge layer indeed correspond to semantically meaningful processing, we visualize the attention patterns as a more interpretable signal of where the model focuses.
-As shown in Figure 7, we visualize attention maps across layers.
-Figure 7: Attention map across layers (GPT-2 Small, Synthetic). At the generalization ridge layers, attention becomes more targeted, focusing on predictive tokens.
-In early layers (e.g., Layer 1), attention is diffuse and biased toward final tokens—reflecting reliance on position rather than true predict signals.
-By generalization ridge layers (e.g., Layer 11), attention becomes more targeted, concentrating on predictive tokens.
-This shift marks a transition from positional attention to semantically meaningful focus, suggesting that intermediate layers are increasingly capable of isolating task-relevant information from irrelevant content.
-In the final layers, attention regresses toward final tokens.
-This reversion is aligned with the observed decline in $I(Z_{\ell};Y)$.
-The resurgence of attention towards terminal tokens indicates a potential memorization scenario, where the model re-engages superficial positional strategies, possibly memorizing noise rather than further refining the generalized extraction of predictive signals.
-This attention trajectory further supports the generalization ridge hypothesis, highlighting a trade-off between generalization and memorization in the model’s representational strategy.
+These scalars are trained separately on the (a) in-distribution (ID) split and (b) out-of-distribution (OOD) split.
+Since no other parameters are updated, the learned $\beta_{\ell}$ serve as a direct diagnostic of the extent to which each layer should be amplified or attenuated to suit the data regime, revealing which layers remain stable across regimes and which adapt strongly to distributional shifts.
+(a) GPT-2 Small on Synthetic Arithmetic
+(b) GPT-2 Small on CLUTRR
+Figure 6: Residual scaling coefficients $\beta_{\ell}$ across all transformer layers.
+ID training emphasizes later layers, while OOD training has lower final layer weight.
+Curve shows the mean across 5 independent run, and the shaded region denotes 1-$\sigma$ error bar.
+As shown in Figure 6, optimizing the residual scaling coefficient parameters on in-distribution data assigns relatively higher residual weights to the final transformer layers, indicating that in-distribution performance benefits from increased reliance on late-layer residual contributions.
+This observation suggests that these deeper layers capture refinements that are closely tied to the training distribution.
+In contrast, under out-of-distribution (OOD) settings, the learned coefficients systematically have relative lower contribution of the final layers compared to the ID settings, revealing that improved generalization is associated with downweighting late-stage residual signals.
+More results are in Appendix E.
+This pattern supports the interpretation that late-layer residuals become increasingly specialized in memorized patterns tied to the training distribution.
+Together, these findings provide empirical support for the generalization ridge hypothesis, demonstrating that information flow in transformers reflects a trade-off between generalizable signals and memorized, distribution-specific features.
 7 Conclusion
-We present InfoRidge, an information-theoretic framework, to analyze how information evolves across layers in transformer language models.
-By estimating both predictive information and incremental information gain, we systematically characterize the dynamics of task-relevant signal propagation throughout the network.
-Our findings reveal a consistent generalization ridge emerging in intermediate layers, where mutual information between the hidden representation and the target label reaches its peak.
+We introduce InfoRidge, an information-theoretic framework designed to trace and quantify how information evolves across layers in transformer-based language models for natural language generation.
+By estimating both predictive information and incremental information gain, we systematically characterize the layerwise dynamics of information flow, offering a principled view of how signals are refined, amplified, or diminished as they propagate through the network.
+Our findings reveal a consistent generalization ridge emerging in intermediate layers, where mutual information between the hidden representation and the target label reaches its peak before gradually declining.
 This phenomenon reflects a fundamental trade-off between generalization and memorization as information flows deeper into the model.
-Residual scaling experiments further corroborate this interpretation, demonstrating the functional specialization of different layers—where intermediate blocks play a key role in supporting generalization, while deeper layers increasingly focus on memorization.
-Attention flow provides additional insight into where the model concentrates information.
-Taken together, these results offer a unified perspective on information flow in transformers for natural language generation and point to opportunities for enhancing model architecture, improving interpretability, and increasing robustness to distributional shifts.
-9 Limitation
-Our analysis focuses on a select set of datasets and model architectures, which may limit the breadth of empirical coverage.
-The mutual information estimates rely on kernel-based approximations that may be sensitive to representation geometry.
-Additionally, the residual scaling approach assumes layer-wise independence and does not account for potential inter-layer dependencies during adaptation.
-Future work could address these limitations by extending the analysis to broader model families and tasks and exploring joint optimization strategies.
+Residual scaling experiments further corroborate this interpretation, demonstrating the functional specialization of different layers.
+Taken together, these findings position InfoRidge as a comprehensive framework for diagnosing how language models internally manage information during natural language generation, while also revealing the structural mechanisms that govern the balance between generalization and memorization in transformer architectures.
+Ethics Statement.
+This work adheres to the Code of Ethics.
+Our experiments use only open-source models and publicly available datasets under their respective open licenses, with no involvement of human subjects or sensitive data.
+We identify no foreseeable ethical risks.
+Reproducibility Statement.
+We ensure reproducibility by providing experimental and implementation details in Section 4 and Appendices BC.
+Full results with statistical significance are in Appendix E, and anonymous source code is included as supplementary material.
 Appendix A Mathematical Details for Matrix-Based Information Estimation and Theoretical Foundations
-We employ the matrix-based Rényi entropy  [11] to estimate mutual information between model representations and target labels.
-This method captured sample similarity structure via kernel Gram matrices.
+We employ the matrix-based Rényi entropy  Giraldo et al. (2014) to estimate mutual information between representations and labels, leveraging kernel Gram matrices to capture sample similarity.
 A.1 Matrix-Based Entropy Estimation
 Let $U=\{u_{i}\}_{i=1}^{N}\subset\mathbb{R}^{d}$ denote $\ell_{2}$-normalized representations obtained from a specific transformer layer, a residual update, or the embedding of the target label.
-A Gaussian kernel Gram matrix $G_{U}\in\mathbb{R}^{N\times N}$ is constructed as:
-$(G_{U})_{ij}=\exp\left(-\frac{\|u_{i}-u_{j}\|^{2}}{2\sigma^{2}}\right),$
-with bandwidth $\sigma=1$. The matrix is then trace-normalized to ensure $\text{tr}(G_{U})=1$.
+A Gaussian kernel Gram matrix $G_{U}\in\mathbb{R}^{N\times N}$ is constructed as: $(G_{U})_{ij}=\exp\left(-\frac{\|u_{i}-u_{j}\|^{2}}{2\sigma^{2}}\right),$ with bandwidth $\sigma=1$.
+The matrix is then trace-normalized to ensure $\text{tr}(G_{U})=1$.
 The matrix-based Rényi entropy of order $\alpha=1$ is defined as:
@@
 A.2 Mutual Information Estimation
-To estimate the mutual information between two random variables $U$ and $V$, we compute their Gram matrices $G_{U}$ and $G_{V}$, and form the joint similarity matrix via element-wise (Hadamard) product:
-$G_{UV}=G_{U}\circ G_{V}.$
-After trace-normalization, mutual information is estimated by:
-$I(U;V)=H(U)+H(V)-H(U,V),$
-where $H(U,V)=-\text{tr}(G_{UV}\log G_{UV})$.
+To estimate the mutual information between two random variables $U$ and $V$, we compute their Gram matrices $G_{U}$ and $G_{V}$, and form the joint similarity matrix via element-wise (Hadamard) product: $G_{UV}=G_{U}\circ G_{V}.$ After trace-normalization, mutual information is estimated by: I(U;V)=H(U)+H(V)H(U,V),I(U;V)=H(U)+H(V)-H(U,V), where $H(U,V)=-\text{tr}(G_{UV}\log G_{UV})$.
 The eigenvalue spectrum of $G_{UV}$ governs the joint entropy term; its shape reflects how much of the structure in $U$ and $V$ aligns.
 A more concentrated spectrum in $G_{UV}$ relative to $G_{U}$ and $G_{V}$ implies stronger dependence and thus higher mutual information.
-A.3 Min Wasserstein Generalization Bound [20]
-We restate the generalization bound proposed by He et al. [20], which characterizes generalization in terms of the Wasserstein distance between internal representations.
-Suppose that the loss function ~$\tilde{\ell}:\mathcal{Y}\times\mathcal{Y}\to\mathbb{R}_{\geq 0}$ is $\rho_{0}$-Lipschitz, and the activation function $\phi_{\ell}:\mathbb{R}\to\mathbb{R}$ is $\rho_{\ell}$-Lipschitz for each $\ell=1,\ldots,L$.
+A.3 Min Wasserstein Generalization Bound (He et al., 2024)
+We restate the generalization bound proposed by He et al. (He et al., 2024), which characterizes generalization in terms of the Wasserstein distance between internal representations.
+Suppose that the loss function $\tilde{\ell}:\mathcal{Y}\times\mathcal{Y}\to\mathbb{R}_{\geq 0}$ is $\rho_{0}$-Lipschitz, and the activation function $\phi_{\ell}:\mathbb{R}\to\mathbb{R}$ is $\rho_{\ell}$-Lipschitz for each $\ell=1,\ldots,L$.
 Then:
-$\small\mathrm{gen}(P_{W|\mathcal{D}_{n}}\!,\!P_{X,Y})\!\leq\!\min_{\ell}\frac{\rho_{0}}{n}\sum_{i=1}^{n}\!\mathbb{E}_{W}\!\left[\!\left(1\!\vee\!\prod_{j=\ell+1}^{L}\!\rho_{j}\|W_{j}\|_{\mathrm{op}}\!\right)\!\cdot\!\mathcal{W}_{1}\!\left(\!P_{T_{\ell},i,Y_{i}|W}(\cdot|W),\!\ P_{T_{\ell},Y|W}(\cdot|W)\!\right)\!\right]$ (3)
-This result shows that generalization can be tightly controlled by the Wasserstein distance [21] between internal representations at a specific layer—referred to as the generalization funnel layer.
+$\mathrm{gen}(P_{W|\mathcal{D}_{n}}\!,\!P_{X,Y})\!\leq\!\min_{\ell}\frac{\rho_{0}}{n}\sum_{i=1}^{n}\!\mathbb{E}_{W}\!\left[\!\left(1\!\vee\!\prod_{j=\ell+1}^{L}\!\rho_{j}\|W_{j}\|_{\mathrm{op}}\!\right)\!\cdot\!\mathcal{W}_{1}\!\left(\!P_{T_{\ell},i,Y_{i}|W}(\cdot|W),\!\ P_{T_{\ell},Y|W}(\cdot|W)\!\right)\!\right]$ (3)
+This result shows that generalization can be tightly controlled by the Wasserstein distance (Villani et al., 2008) between representations at a specific layer—referred to the generalization funnel layer.
 Connection to Our Work.
-The Min Wasserstein Generalization Bound [20] provides a theoretical foundation for our study by characterizing generalization in terms of distributional alignment at an intermediate layer.
+The Min Wasserstein Generalization Bound (He et al., 2024) provides a theoretical foundation for our study by characterizing generalization in terms of distributional alignment at an intermediate layer.
 It motivates our analysis of information flow by suggesting that information is peaked at a specific layer—the generalization funnel.
-Our InfoRidge framework builds on this insight by quantifying predictive information across layers, and reveals that a specific intermediate layer exhibits peak mutual information and correlates with better generalization performance.
+Our InfoRidge builds on this insight by quantifying predictive information across layers, and reveals that a specific intermediate layer exhibits peak mutual information and correlates with better generalization performance.
 Appendix B Dataset Overview and Statistics
@@
 B.2 CLUTRR
-CLUTRR (Compositional Language Understanding and Text-based Relational Reasoning) [25] is a diagnostic benchmark for evaluating relational inference in language models.
+CLUTRR (Compositional Language Understanding and Text-based Relational Reasoning) (Sinha et al., 2019) is a diagnostic benchmark for evaluating relational inference in language models.
 Each example contains a story describing family relations, and the task is to infer the missing relationship between two entities.
@@
 B.3 ECQA
-ECQA (Explanations for CommonsenseQA) [26] is a commonsense multiple-choice question-answering dataset, where each question is accompanied by 5 answer options.
+ECQA (Explanations for CommonsenseQA) (Aggarwal et al., 2021) is a commonsense multiple-choice question-answering dataset, where each question is accompanied by 5 answer options.
 B.4 Synthetic Arithmetic Dataset
@@
 Each element in the sequence is represented as a string of the form:
-[…]
+$\texttt{S\{signal\}\_N\{noise\}},\quad\text{where noise}\sim\mathcal{U}_{\text{int}}(0,\texttt{noise\_range-1}).$
 Here $\mathcal{U}_{\text{int}}$ denotes the uniform distribution. The model is trained to predict the signal value of the final (10th) element, using the preceding elements as input context.
@@
 C.1 Prompt Construction
-All tasks are cast into a next-token generation format.
-The model receives a prompt (context) and must generate the correct continuation token.
-Below are construction strategies and examples for each dataset:
+All tasks are cast into a next-token generation format. The model receives a prompt and generate the next token. Below are construction strategies and examples for each dataset:
 CLUTRR
@@
 This controlled format enables manipulation of distributional properties by varying the modulus $K$.
-C.2 Information Estimation
-To estimate mutual information, we subsample between 50 and 200 test examples, with the exact number determined by the model and task, in order to avoid excessive memory usage from large Gram matrices.
-Full results and statistical significance are reported in Appendix E.
-C.3 Residual Scaling with Learnable $\beta_{\ell}$ Parameters
+C.2 Finetuning Settings
+We fine-tuned all layers end-to-end using AdamW (learning rate $5\times 10^{-6}$, weight decay $0.01$) with a linear schedule (warmup ratio $0.1$) and early stopping on validation loss.
+Training converged in all settings.
+C.3 Information Estimation
+To estimate mutual information, we subsample between 50 and 200 test examples depending on model and task to achieve a stable result.
+C.4 Residual Scaling with Learnable $\beta_{\ell}$ Parameters
 We introduce a vector of learnable scalar weights $\beta=\{\beta_{1},\dots,\beta_{L}\}$ applied to residual connections in a frozen transformer:
@@
   • Each transformer block’s residual output is modulated by $\beta_{\ell}$ via forward hooks.
-  • $\beta$ values are initialized to 1 and updated using gradient descent without altering model architecture.
+  • $\beta$ are initialized to 1 and updated using gradient descent without altering the architecture.
 Training and Evaluation Settings
+Table 4: Performance (%) of different models across datasets.
+Model CLUTRR Synthetic ECQA
+GPT-2 Small 30.98 71.45
+GPT-2 Medium 34.21 68.35
+Qwen-2.5-0.5B 31.94 83.20 57.06
+LLaMA-3.1-8B 57.68 94.00 77.39
 We perform residual $\beta_{\ell}$ training on top of a model that has already been fine-tuned on the target dataset, keeping all previously learned weights frozen and optimizing only the $\beta_{\ell}$ parameters using HuggingFace’s Trainer API.
@@
 Computing Resources
-All experiments were conducted on an NVIDIA RTX A6000 GPU with 48GB of memory.
+Experiments were conducted on an NVIDIA RTX A6000 GPU (48GB).
 Model Licenses
-The GPT-2 Small [22] and GPT-2 Medium [22] models are released under the MIT License and are available via Hugging Face Transformers.
-The Qwen-2.5 0.5B [23] model is provided by Alibaba Group under the Apache 2.0 License.
-The LLaMA-3.1 8B [24] model is made available by Meta under a non-commercial research license and accessed via Hugging Face.
+The GPT-2 Small (Radford et al., 2019) and GPT-2 Medium (Radford et al., 2019) models are released under the MIT License and are available via Hugging Face Transformers.
+The Qwen-2.5 0.5B (Yang et al., 2024) model is provided by Alibaba Group under the Apache 2.0 License.
+The LLaMA-3.1 8B (Meta AI, 2024) model is made available by Meta under a non-commercial research license and accessed via Hugging Face.
 Dataset Licenses
-CLUTRR [25] is released under a CC BY 4.0 license as part of EMNLP 2019.
-The ECQA [26] dataset is also released under a CC BY 4.0 license by its authors as part of EMNLP 2021.
+CLUTRR (Sinha et al., 2019) is released under a CC BY 4.0 license as part of EMNLP 2019.
+The ECQA (Aggarwal et al., 2021) dataset is also released under a CC BY 4.0 license by its authors as part of EMNLP 2021.
 The Synthetic Arithmetic dataset is custom-designed by the authors and does not rely on any external or licensed data sources.
@@
 Appendix E Full Quantitative Results with Confidence Estimates
-To provide a comprehensive view of model behavior across layers, we report the full experimental results for all models and datasets evaluated in this study.
-This includes predictive information trends ($I(Z;Y)$), incremental information gain ($I(\Delta Z;Y)$) and residual scaling effects ($\beta_{\ell}$), together with statistical significance reporting.
+To provide a comprehensive view of model behavior, we report the full experimental results for all models and datasets.
+This includes predictive information trends ($I(Z;Y)$), incremental information gain ($I(\Delta Z;Y)$) and residual scaling effects ($\beta_{\ell}$), with statistical significance reporting.
 E.1 Predictive Information
-To estimate mutual information between hidden representations and target labels, we compute matrix-based mutual information following the methodology outlined in Appendix A.
-Due to the high memory cost of computing $N\times N$ Gram matrices, we subsample between 50 and 100 test examples per experiment and report results with statistical significance based on multiple random seeds.
+To estimate mutual information between hidden representations and target labels, we compute matrix-based mutual information following the methodology in Appendix A.
+We subsample between 50 and 100 test examples per experiment and report results with statistical significance based on multiple random seeds.
 (a) GPT-2 Small on CLUTRR
@@
 Lighter line colors represent later training epochs.
-Each curve shows the mean across five random seeds (0, 1, 2, 3, 42), and the shaded region denotes a 2-sigma (~96%) confidence interval.
+Each curve shows the mean across 5 random seeds (0, 1, 2, 3, 42), and the shaded region denotes a 2-sigma (~96%) confidence interval.
 E.2 Incremental Information Gain
 In addition to predictive information $I(Z^{(\ell)};Y)$, we compute the incremental information gain $I(\Delta Z^{(\ell)};Y)$ at each layer.
-This quantity the information contribution for each transformer blocks.
+This quantifies the information contribution for each transformer block.
 (a) GPT-2 Small on CLUTRR
@@
 E.3 Residual Scaling
-We introduce a residual scaling mechanism with learnable scalar coefficient parameters, inspired by prior work on adaptive residual modulation [12, 13].
-Similar ideas of modulating internal computation have also been explored in parameter-efficient fine-tuning (PEFT) [27], Representation-Efficient Fine-Tuning (REFT) [28] and interpretability-driven control [29, 30, 31, 32].
+We introduce a residual scaling mechanism with learnable scalar coefficient parameters, inspired by prior work on adaptive residual modulation (Liu et al., 2019a; Menghani et al., 2024).
+Similar ideas of modulating internal computation have also been explored in parameter-efficient fine-tuning (PEFT) (Houlsby et al., 2019), Representation-Efficient Fine-Tuning (REFT) (Wu et al., 2024) and interpretability-driven control (Huang et al., 2024; Deng et al., 2025; Meng et al., 2022; Wu et al., 2023).
 We present the complete set of residual scaling results, detailing the learned $\beta_{\ell}$ values across all transformer layers.
@@
 We observe that in the ID setting, later layers tend to receive higher weights, consistent with memorization behavior.
-In contrast, OOD training consistently downweights final layers and shifts importance toward the middle of the network, aligning with the generalization ridge identified through InfoRidge.
-(a) GPT-2 Small on Synthetic Arithmetic
-(b) GPT-2 Small on CLUTRR
-(c) Qwen-2.5-0.5B on Synthetic Arithmetic
-(d) Qwen-2.5-0.5B on CLUTRR
-(e) LLaMA-3.1-8B on Synthetic Arithmetic
-(f) LLaMA-3.1-8B on CLUTRR
+In contrast, OOD training consistently has lower final layers’ weight compared to the ID settings and shifts importance toward the intermediate layers.
+When the task is simple relative to the model’s capability, such as LLaMA on the synthetic task, the final-layer residual coefficients are similar for ID and OOD, consistent with Figure 7.
+The in-distribution (ID) setting is using the training split of “gen_train23_test2to10”.
+Out-of-distribution (OOD) is conducted on held-out CLUTRR configurations that do not overlap with the training distribution: for GPT models we evaluate on the test split of “gen_train23_test2to10” (lengths 2 through 10), while for Qwen and LLaMA we use the test split of “rob_train_sup_23_test_all_23”, which defines a more challenging OOD setting by jointly varying both the clause length and the status of noise facts introduced into the story, given the higher modeling capacity of these models.
+In all cases, tasks 1.2 and 1.3 are excluded from the OOD sets, as they overlap with the training configuration and are therefore treated as in-distribution rather than out-of-distribution.
+(a) GPT-2 Small on CLUTRR
+(b) GPT-2 Small on Synthetic Arithmetic
+(c) Qwen-2.5-0.5B on CLUTRR
+(d) Qwen-2.5-0.5B on Synthetic Arithmetic
+(e) LLaMA-3.1-8B on CLUTRR
+(f) LLaMA-3.1-8B on Synthetic Arithmetic
 Figure 9: Residual scaling coefficients $\beta_{\ell}$ across all transformer layers.
-ID training emphasizes later layers, while OOD training shifts weight toward middle layers, aligning with the generalization ridge observed via InfoRidge.
+ID training emphasizes later layers, while OOD training has lower final layer weight.
 Each curve shows the mean across five random seeds (0, 1, 2, 3, 42), and the shaded region denotes 1-sigma error bar.
-Appendix F Incremental Information Gain Case Study
+Appendix F Attention Dynamics Across Layers Case Study
+To verify whether generalization ridge layer indeed correspond to semantically meaningful processing, we visualize the attention patterns as a more interpretable signal of where the model focuses.
+As shown in Figure 10, we visualize attention maps across layers.
+Figure 10: Attention map across layers (GPT-2 Small, Synthetic). At the generalization ridge layers, attention becomes more targeted, focusing on predictive tokens.
+In early layers (e.g., Layer 1), attention is diffuse and biased toward final tokens—reflecting reliance on position rather than true predict signals.
+By generalization ridge layers (e.g., Layer 11), attention becomes more targeted, concentrating on predictive tokens.
+This shift marks a transition from positional attention to semantically meaningful focus, suggesting that intermediate layers are increasingly capable of isolating task-relevant information from irrelevant content.
+In the final layers, attention regresses toward final tokens.
+This reversion is aligned with the observed decline in $I(Z_{\ell};Y)$.
+The resurgence of attention towards terminal tokens indicates a potential memorization scenario, where the model re-engages superficial positional strategies, possibly memorizing noise rather than further refining the generalized extraction of predictive signals.
+This attention trajectory further supports the generalization ridge hypothesis, highlighting a trade-off between generalization and memorization in the model’s representational strategy.
+Appendix G Incremental Information Gain Case Study
 To further illustrate how intermediate layers contribute to generalization, we analyze the semantic content introduced by residual transitions at different depths.
@@
 This analysis allows us to inspect the linguistic shift introduced by each transformer block in isolation, and to assess whether the changes correspond to task-relevant predictions or superficial noises.
-F.1 Methodology
+G.1 Methodology
 For a fixed input, we compute residual transitions $\delta z^{(\ell)}$ at each layer and pass them through the model’s final linear projection (LM head) followed by a softmax.
 We record the top predicted tokens and their probabilities, and it reflects the directional change applied by layer $\ell$.
-F.2 Observations
+G.2 Observations
 We decode residual transitions across layers and report the top shifted tokens by projecting $\Delta Z^{(\ell)}$ through the LM head.
@@
 These patterns are consistent with our mutual information analysis, which identifies intermediate layers as better semantically aligned with the prediction target—forming a ridge of generalization.
-F.3 Implication
+G.3 Implication
 These decoding results reinforce our interpretation of the generalization ridge: intermediate layers contribute the most semantically informative updates to the model’s representation.
 The residual transitions thus serve as a useful lens for understanding how and where semantic meaning is introduced during forward propagation.
+Appendix H Activation Patching Experiment
+To assess which residual blocks causally support OOD prediction, we conduct a targeted activation-patching experiment.
+Let the model contain $L$ residual blocks with hidden states $Z_{\ell}\in\mathbb{R}^{d}$, and let $t^{*}$ denote the final context token (the last token before next-token prediction).
+For every OOD sequence $x_{\mathrm{ood}}$, we pair it with an ID sequence $x_{\mathrm{id}}$ sharing the same token structure.
+For each depth $\ell\in\{1,\dots,L\}$, we compute two forward passes: a standard OOD pass and a patched pass in which we replace only the hidden state of token $t^{*}$ at depth $\ell$ according to $Z_{\ell}(x_{\mathrm{ood}},t^{*})\leftarrow Z_{\ell}(x_{\mathrm{id}},t^{*}).$ All other activations remain unchanged.
+Let $logit_{\mathrm{base}}$ denote the logit assigned to the correct next token under the baseline OOD pass, and let $logit_{\mathrm{patch}}^{(\ell)}$ denote the corresponding logit under layer-$\ell$ patching.
+We measure causal impact using the logit drop $\Delta_{\ell}=logit_{\mathrm{base}}-logit_{\mathrm{patch}}^{(\ell)},$ where a positive value indicates that depth $\ell$ provides information necessary for correct OOD prediction.
+Table 5 reports the averaged logit drop for all $\ell=1,\dots,12$ (GPT-2 Small on Synthetic Dataset).
+The effects are negligible for layers 1–8, begin rising around layer 9, increase sharply at layer 10, peak at layer 11, and decline again at layer 12.
+This rise–peak–fall pattern closely matches the ridge structure observed in Figure 2, where $I(Z_{\ell};Y)$ increases through mid-to-late depth, reaches a maximum in the ridge region, and then decreases.
+The activation-patching experiment therefore provides supporting causal evidence that the ridge layers identified by IndoRidge are necessary for successful OOD generalization.
+Table 5: Average logit drop $\Delta_{\ell}$ (baseline logit - patched logit). Larger values indicate greater causal contribution to OOD prediction.
+Layer $\ell$ 1 2 3 4 5 6 7 8 9 10 11 12
+$\Delta_{\ell}$ -0.1037 -0.1039 -0.0327 -0.5405 -5.1663 -8.3992 -5.8914 -15.0996 -3.0998 9.3148 17.2022 15.5943
+Appendix I Extending Predictive Information to Multi-Token Outputs
+To complement our single-token analysis, we further evaluate information flow under multi-token outputs using the CNN/Daily Mail summarization dataset (1.0.0, test) (See et al., 2017; Hermann et al., 2015).
+Autoregressive next-token information.
+We follow the model’s natural decoding process.
+We analyze $I(Z_{\ell}^{(t)};\hat{Y}^{(t+1)})$ since there is no deterministic gold answer for each intermediate step in NLG, where for each generation step $t$, $Z_{\ell}^{(t)}$ is the hidden state of current token at layer $\ell$ and $\hat{Y}^{(t+1)}$ is the embedding of model’s predicted token at step $t+1$.
+Across steps, the pattern is consistent (Figure 11): information rises from early layers, peaks in the middle layers, and then declines toward top layers.
+This gives a clear information ridge, indicating that intermediate layers encode the most stable and task-relevant core information that supports the model’s decision-making (generalization).
+In particular, high $I(Z_{\ell}^{(t)};\hat{Y}^{(t+1)})$ at these layers means the model’s predictive direction is most strongly determined by that layer: layers whose representations consistently constrain the next-token prediction across many contexts must encode the shared, invariant structure that the model relies on to generalize.
+In contrast, upper layers exhibit behavior that might be related to pretrained surface patterns from corpus such as token co-occurrence, causing the representation to drift away from this decision-relevant structure (memorization).
+The first step is a notable outlier: it is the only step that conditions solely on the input prompt, and shows the sharpest and most pronounced ridge.
+Generating the first token requires a global transition from the prompt into the answer trajectory, which heavily engages the abstract, decision-relevant subspace in the middle layers.
+Crucially, this is also the point where information compression is most extreme, the model must condense all prompt-level semantics into a single initial token decision, therefore forms a substantial ridge.
+Once generation begins, later generation steps move into a more regular autoregressive regime, where each hidden state already contains a compressed summary of all previously generated content, and the continuation depends more on local context than on global restructuring.
+Consequently, the depth profiles become flatter and more stable across steps, although the mid-layer ridge remains visible.
+(a) Qwen-2.5-0.5B on CNN/Daily Mail for 50 generation steps.
+(b) LLaMA-3.1-8B on CNN/Daily Mail for 50 generation steps.
+Figure 11: $I(Z_{\ell}^{(t)};\hat{Y}^{(t+1)})$ across layers (x-axis) and generation steps (y-axis), where the left y-axis shows first step mutual information (bars) and the right y-axis shows mutual information for the rest steps.
+Pooled-answer information.
+We quantify how much information each layer carries about the gold answer by computing $I(Z_{\ell}^{(t)};Y^{\text{pool}})$, where $Y^{\text{pool}}$ is a fixed target vector obtained by average-pooling the embedding-layer representations of all ground-truth answer tokens.
+Across models and steps (Figure 12), a consistent pattern emerges: intermediate layers are where the hidden state is most strongly aligned with the gold-answer.
+(a) Qwen-2.5-0.5B on CNN/Daily Mail for 50 generation steps.
+(b) LLaMA-3.1-8B on CNN/Daily Mail for 50 generation steps.
+Figure 12: $I(Z_{\ell}^{(t)};Y^{\text{pool}})$ across layers (x-axis) and generation steps (y-axis), where the left y-axis shows first step mutual information (bars) and the right y-axis shows mutual information for the rest steps.
+Taken together, the two constructions reveal a coherent picture of depth specialization in multi-token generation.
+Intermediate layers are simultaneously the most informative about the model’s next-token decisions and the most aligned with the global answer.
+Thus, the two observed ridges are complementary projections of a single underlying phenomenon: intermediate layers concentrated most meaningful and decision-relevant information.
+In contrast, upper layers increasingly reflect pretrained corpus-specific surface patterns, causing the representation to drift away from this decision-relevant structure.
+(a) $I(Z_{\ell}^{(t)};\hat{Y}^{(t+1)})$: Qwen-2.5-0.5B on CNN/Daily Mail for 50 generation steps.
+(b) $I(Z_{\ell}^{(t)};\hat{Y}^{(t+1)})$: LLaMA-3.1-8B on CNN/Daily Mail for 50 generation steps.
+Figure 13: Mutual information comparing stopwords/punctuation against content words, showing that both stopwords/punctuation have consistently higher MI.
+We also plot the mutual information comparing stopwords/punctuation against content words, showing that both stopwords/punctuation have consistently higher MI (Figure 13).
+These tokens are more predictable because they follow strict grammatical and structural patterns, whereas content words depend more on semantic context and therefore exhibit greater variability.
+Appendix J Non-finetuned scenario
+Beyond fine-tuning scenarios, we also computed predictive-information $I(Z_{\ell};Y)$ curves in a non-training zero-shot regime and observed the same mid-layer information peak pattern, see Figure 14.
+Figure 14: GSM8K on Qwen2.5-0.5B shows an intermediate-layer MI rise and late-layer decline.
+Appendix K Kernel Ablation
+The matrix-based MI estimator requires a positive-definite Gram matrix, for which the Gaussian kernel is the standard choice.
+To assess robustness with respect to kernel selection, we additionally test the Laplacian and Polynomial kernels.
+While these kernels introduce quantitative shifts in the magnitude, the trend—including the rise–peak–decline structure and the location of the ridge—remains unchanged (Figure 15).
+Polynomial Kernel.
+Given a set of normalized representations $U=\{u_{i}\}_{i=1}^{N}$, the polynomial kernel Gram matrix is defined as
+$(G^{\mathrm{poly}}_{U})_{ij}=\left(u_{i}^{\top}u_{j}+c_{0}\right)^{p},$ (4)
+where $p$ is the polynomial degree and $c_{0}$ is a constant bias term.
+Laplacian Kernel.
+Given the same representation set $U=\{u_{i}\}_{i=1}^{N}$, the Laplacian kernel Gram matrix is defined as
+$(G^{\mathrm{lap}}_{U})_{ij}=\exp\!\left(-\gamma\,\|u_{i}-u_{j}\|_{1}\right),$ (5)
+where $\gamma>0$ is a kernel coefficient and $\|\cdot\|_{1}$ denotes the $\ell_{1}$ distance.
+(a) Gaussian Kernel: Qwen-2.5-0.5B on ECQA.
+(b) Polynomial Kernel: Qwen-2.5-0.5B on ECQA.
+(c) Laplacian Kernel: Qwen-2.5-0.5B on ECQA.
+Figure 15: Different kernel shows similar pattern.
+Appendix L Use of Large Language Models.
+Large language models (LLMs) were used solely as assistive tools for proofreading and improving clarity of writing.

### v3

@@
 While prior studies suggest that intermediate layers often yield more generalizable representations than final layers, how this generalization ability emerges and propagates across layers during training remains unclear.
-To address this gap, we propose InfoRidge, an information-theoretic framework, to characterize how predictive information—the mutual information between hidden representations and target outputs—varies across depth.
-Estimating this quantity enables us to trace the flow of task-relevant information throughout the model during training.
-Our experiments across various models and datasets reveal a consistent non-monotonic trend: predictive information peaks in upper-middle layers—forming a generalization ridge—before declining in final layers, reflecting a transition between generalization and memorization.
-To further investigate this phenomenon, we introduce residual scaling coefficients—trainable scalar parameters applied to each residual block—which serve as functional probes for assessing the relative importance of individual transformer layers.
-These coefficients reveal that, under distribution shift, models downweight final layers and increasingly rely on intermediate layers, highlighting their role in generalization.
+To address this gap, we propose InfoRidge, an information-theoretic framework, to characterize how predictive information—the mutual information between hidden representations and target outputs—varies across depth during training.
+Our experiments across various models and datasets reveal a consistent non-monotonic trend: predictive information peaks in intermediate layers—forming a generalization ridge—before declining in final layers, reflecting a transition between generalization and memorization.
+To further investigate this phenomenon, we conduct a set of complementary analyses that leverage residual scaling, attention pattern, and controlled model capacity to characterize layer-wise functional specialization.
+We further validate our findings with multiple-token generation experiments, verifying that the observed ridge phenomenon persists across decoding steps.
 Together, these findings offer new insights into the internal mechanisms of transformers and underscore the critical role of intermediate layers in supporting generalization.
@@
 Nevertheless, we lack a rigorous understanding of how these models acquire and synthesize task-relevant information during training.
-A growing body of research has shown that intermediate layers in deep neural networks often surpass final layers in terms of representational quality and generalization performance (Liu et al., 2019b; Voita et al., 2019; Ansuini et al., 2019; Ahrens et al., 2023; Uselis & Oh, 2025).
+A growing body of research has shown that intermediate layers in deep neural networks often surpass final layers in terms of representational quality and generalization performance (Liu et al., 2019b; Voita et al., 2019; Ansuini et al., 2019; Ahrens et al., 2023; Uselis and Oh, 2025).
 In language models, intermediate layers often encode richer semantic and more robust features than final layers (Fan et al., 2024; Jin et al., 2024; Skean et al., 2025).
@@
 This analysis directly connects the information peak to generalizable behavior and clarifies how generalization and memorization are distributed across depth.
-To further validate this interpretation, we introduce residual scaling coefficients—learnable scalar parameters $\beta_{\ell}$ applied to each residual block—while keeping all other model weights frozen, drawing inspiration from prior work on layer-wise adaptation (Liu et al., 2019a; Menghani et al., 2024).
-A higher $\beta_{\ell}$ value indicates greater reliance on the corresponding layer’s output during prediction.
-These coefficients act as functional probes, revealing how the model redistributes layer-wise importance under different data distribution.
-Under in-distribution training, deeper layers retain higher residual weights, reflecting the model’s reliance on memorized, task-specific features.
-When evaluated under distribution shift, models have lower reliance on the final layer and shifts importance toward the intermediate layers.
-To understand the formation of the generalization ridge, we analyze both attention patterns and model capacity.
+We further extend InfoRidge to multi-token natural language generation.
+By treating the full output sequence as a structured target and measuring mutual information between hidden states and the long-form target distribution, we show that the generalization ridge persists across decoding steps.
+Despite increasing mutual information as generation progresses, the layer-wise structure remains stable: predictive information consistently peaks in early-to-intermediate layers and declines in later layers.
+To understand the formation of the generalization ridge, we analyze attention patterns, model capacity and residual scaling coefficients.
 Attention analysis shows ridge layers attend to tokens that capture broadly useful features, aligning with the information peak.
-Beyond attention patterns, we also investigate the conditions under which the ridge emerges.
+We also investigate the conditions under which the ridge emerges.
 Our depth ablation results show that the ridge only emerges beyond a certain depth threshold.
 Below this threshold, predictive information increases monotonically, indicating that sufficient capacity is a prerequisite for generalization ridge to emerge.
+We also introduce residual scaling coefficients $\beta_{\ell}$—learnable scalars applied to each residual block with all other weights frozen—as functional probes of layer-wise reliance (Liu et al., 2019a; Menghani et al., 2024).
+When evaluated under distribution shift, models have lower reliance on the final layer and shifts importance toward the intermediate layers.
 Contributions.
@@
 2. We introduce InfoRidge, an information-theoretic framework that applies matrix-based mutual information estimation to autoregressive language models to analyze information flow.
-3.
-We introduce residual scaling coefficients as trainable indicators of how models shift representational focus across layers during training, providing a causal, adaptive measure of model behavior under distribution shifts.
+3. We analyze attention patterns, model capacity and residual scaling coefficients, providing causal evidences for the generalization ridge phenomenon.
 2 Related Work
-Understanding how information is encoded and transformed across layers has been studied through probing classifiers (Alain & Bengio, 2016), attention flow (Vig & Belinkov, 2019), and information-theoretic approaches such as the information bottleneck (Shwartz-Ziv & Tishby, 2017), mutual information estimation (Goldfeld, 2019), and matrix-based entropy (Giraldo et al., 2014), offering different lenses to quantify representational capacity, abstraction, and invariance across layers.
-A growing body of research has shown that intermediate layers in deep networks often outperform final layers in terms of representational quality and task performance (Ansuini et al., 2019; Yosinski et al., 2014; Uselis & Oh, 2025; Ahrens et al., 2023; Ando et al., 2023).
+Understanding how information is encoded and transformed across layers has been studied through probing classifiers (Alain and Bengio, 2016), attention flow (Vig and Belinkov, 2019), and information-theoretic approaches such as the information bottleneck (Shwartz-Ziv and Tishby, 2017), mutual information estimation (Goldfeld, 2019), and matrix-based entropy (Giraldo et al., 2014), offering different lenses to quantify representational capacity, abstraction, and invariance across layers.
+A growing body of research has shown that intermediate layers in deep networks often outperform final layers in terms of representational quality and task performance (Ansuini et al., 2019; Yosinski et al., 2014; Uselis and Oh, 2025; Ahrens et al., 2023; Ando et al., 2023).
 In language models, mid-depth layers tend to capture richer semantic or robust features than output layers (Liu et al., 2019b; Voita et al., 2019; Jin et al., 2024; Fan et al., 2024).
@@
 These findings challenge the assumption that deeper layers always yield better representations.
-This pattern holds across settings such as transfer learning (Yosinski et al., 2014), continual learning (Ahrens et al., 2023), and out-of-distribution generalization (Uselis & Oh, 2025).
+This pattern holds across settings such as transfer learning (Yosinski et al., 2014), continual learning (Ahrens et al., 2023), and out-of-distribution generalization (Uselis and Oh, 2025).
 Furthermore, recent work has evaluated representation quality using entropy, curvature, and invariance (Skean et al., 2025), while other studies have examined embedding drift and representational geometry (Merchant et al., 2020; Dar et al., 2022), analyzed memorization and factual recall (Haviv et al., 2022; Yu et al., 2023), and introduced causal perspectives on layer importance through mediation analysis and targeted interventions (Vig et al., 2020; Meng et al., 2022).
 Training-dynamics studies investigated how earlier models develop and refine semantic features across depth, providing an additional perspective on the evolution of layer-wise representations (Merchant et al., 2020; Kumar et al., 2023).
-Correlational probes (e.g., linear probes (Alain & Bengio, 2016)) measure only whether a feature can be decoded from a representation, which reflects correlation but not causal influence.
+Correlational probes (e.g., linear probes (Alain and Bengio, 2016)) measure only whether a feature can be decoded from a representation, which reflects correlation but not causal influence.
 In contrast, causal methods intervene on internal activations to test how changing a component alters the model’s prediction, thereby identifying true causal contribution rather than mere feature presence.
-Our residual-scaling approach aligns with this causal perspective at the layer level.
+Our residual-scaling approach aligns with this causal perspective at layer level.
 However, the underlying causes and functional role of this phenomenon remain only partially understood, motivating further investigation.
@@
 We reveal a consistent non-monotonic peak in the middle layers—termed the generalization ridge—that reflects a meaningful transition in representational focus and aligns with stronger generalization behavior.
-Additionally, unlike prior work focused on classification tasks, we extend the analysis of information flow to generation tasks by quantifying the mutual information between hidden states and the next token.
-This enables us to understand how generalization and memorization dynamics evolve for next-token generation setting during training from an information-theoretic perspective.
+Additionally, unlike prior work focused on classification tasks, we extend the analysis of information flow to generation tasks, enabling us to understand how generalization and memorization dynamics evolve for natural language generation setting from an information-theoretic perspective.
 3 InfoRidge: Information Estimation Framework
@@
 Across a batch of $N$ sequences, we collect the representations:
-$Z_{\ell}=\{z^{(\ell)}_{i}\}_{i=1}^{N},\quad Y=\{y_{i}\}_{i=1}^{N},\quad\Delta Z_{\ell}=\{\delta z_{i}^{(\ell)}\}_{i=1}^{N},$
+$Z_{\ell}=\{z^{(\ell)}_{i}\}_{i=1}^{N}\!,\;Y=\{y_{i}\}_{i=1}^{N}\!,\;\Delta Z_{\ell}=\{\delta z_{i}^{(\ell)}\}_{i=1}^{N},$
 where all vectors are $\ell_{2}$-normalized.
@@
 A positive-definite Gram matrix $G_{U}\in\mathbb{R}^{N\times N}$ is computed using a Gaussian kernel $\kappa$ with bandwidth set to 1 and the matrix is then trace-normalized to satisfy $\operatorname{tr}(G_{U})=1$.
+To assess robustness with respect to kernel selection, we additionally test the Laplacian and Polynomial kernels.
+While these kernels introduce quantitative shifts in the magnitude, the trend remains unchanged, see Appendix E.
 The matrix-based Rényi entropy (with order $\alpha=1$) is then given by:
@@
 The mutual information between them is computed as:
-$I(\mathcal{U};\mathcal{V})\approx I(G_{\mathcal{U}};G_{\mathcal{V}})=H(G_{\mathcal{U}})+H(G_{\mathcal{V}})-H(G_{\mathcal{U}}\circ G_{\mathcal{V}}),$ (2)
-where $\circ$” denotes the Hadamard (elementwise) product. Mathematical details are in Appendix A.
+$I(\mathcal{U};\mathcal{V})$ $\approx I(G_{\mathcal{U}};G_{\mathcal{V}})$ (2)
+$=H(G_{\mathcal{U}})+H(G_{\mathcal{V}})-H(G_{\mathcal{U}}\circ G_{\mathcal{V}}),$
+where $\circ$” denotes the Hadamard product. Mathematical details are in Appendix A.
 4 Experimental Setup
@@
 Datasets.
-We assess model behavior across three tasks casted into NLG problems: CLUTRR (Sinha et al., 2019), a relational reasoning benchmark; ECQA (Aggarwal et al., 2021), a commonsense QA benchmark; and Synthetic Arithmetic, a controlled dataset designed to disentangle task-relevant signal from noise.
+We assess model behavior across three tasks casted into NLG problems: CLUTRR (Sinha et al., 2019), a relational reasoning benchmark; ECQA (Aggarwal et al., 2021), a commonsense QA benchmark; CNN/DailyMail (See et al., 2017; Hermann et al., 2015), a summarization benchmark and Synthetic Arithmetic, a controlled dataset designed to disentangle task-relevant signal from noise.
 Dataset and implementation details are in Appendix B and C.
@@
 Across models and tasks, predictive information curves exhibit a consistent three-phase pattern:
-Table 1: Information Dynamics and Layer-wise Performance (GPT-2-S, Synthetic). OOD performance declines beyond the generalization ridge.
-Layer $\mathbf{I(Z;Y)}$ Test Accuracy (%) All In-Dist.
-Out-Dist.
-Layer 1 0.0782 0.00 0.00 0.00 Layer 2 0.0776 0.00 0.00 0.00 Layer 3 0.0907 0.00 0.00 0.00 Layer 4 0.1156 0.00 0.00 0.00 Layer 5 0.1222 0.00 0.00 0.00 Layer 6 0.1275 0.00 0.00 0.00 Layer 7 0.1552 0.75 0.00 1.50 Layer 8 0.2594 19.45 0.00 40.70 Layer 9 0.2937 20.60 0.60 42.00 Layer 10 0.3295 64.25 72.90 56.90 Layer 11 0.3296 71.65 100.00 42.90 Layer 12 0.1501 70.45 100.00 41.40
 Progressive Accrual (early layers).
@@
 These layers appear to play a critical role in synthesizing abstract features that are essential for generalization.
+Table 1: Information Dynamics and Layer-wise Performance (GPT-2-S, Synthetic). OOD performance declines beyond the generalization ridge.
+Layer $\mathbf{I(Z;Y)}$ Test Accuracy (%) All In-Dist.
+Out-Dist.
+Layer 1 0.0782 0.00 0.00 0.00 Layer 2 0.0776 0.00 0.00 0.00 Layer 3 0.0907 0.00 0.00 0.00 Layer 4 0.1156 0.00 0.00 0.00 Layer 5 0.1222 0.00 0.00 0.00 Layer 6 0.1275 0.00 0.00 0.00 Layer 7 0.1552 0.75 0.00 1.50 Layer 8 0.2594 19.45 0.00 40.70 Layer 9 0.2937 20.60 0.60 42.00 Layer 10 0.3295 64.25 72.90 56.90 Layer 11 0.3296 71.65 100.00 42.90 Layer 12 0.1501 70.45 100.00 41.40
 Representational Compression (final layers).
@@
 Overfitting Adds Memorization in Final Layers
-Figure 4: $I(Z_{\ell};Y)$ rises in final layers during overfitting (LLaMA, ECQA).
 To probe overfitting dynamics, we intentionally fine-tuned the model beyond the optimal point.
@@
 After overfitting, however, the later-layer rise reflects a shift toward memorization, where the model begins to encode superficial shortcuts or redundant label-specific noise rather than useful task information.
+Figure 4: $I(Z_{\ell};Y)$ rises in final layers during overfitting (LLaMA, ECQA).
 Semantically Important Attention Peaks Where Information Peaks
@@
 For example, in Qwen-2.5-0.5B on ECQA, signal attention peaks at Layer 17 (0.2458) but drops to 0.0066 in the final layer, where last-token attention dominates (0.1697).
-Additional results are in Appendix F.
 Table 2: Average attention statistics: (1) average attention scores over all tokens, (2) average attention to signal tokens, (3) the maximum signal attention and its corresponding layer, (4) signal token attention in the final layer and (5) last token attention in the final layer.
@@
 This concentration of information gain further underscores their central role in encoding those task-relevant features that are essential for supporting generalization.
-For additional illustration, a detailed case study that decodes $\Delta z$ via the LM head is provided in Appendix G, highlighting fine-grained token-level shifts and variations observed across layers.
-Figure 5: Middle blocks are key to encoding generalizable information (GPT-2-S, Synthetic).
 The Incremental Information Gain analysis reveals a clear pattern: middle transformer blocks are key to encoding generalizable task-relevant information, thereby forming a generalization ridge.
@@
 This trend underscores a fundamental trade-off in transformer training dynamics.
-6 Residual Scaling Dynamics via Learnable $\beta$ Coefficients
+Additional results are in Appendix E.
+Figure 5: Middle blocks are key to encoding generalizable information (GPT-2-S, Synthetic).
+Residual Scaling Dynamics via Learnable $\beta$ Coefficients
 To deepen our understanding of the Generalization Ridge hypothesis, we examine how modulating the contribution of individual transformer blocks affects information flow, and propose a corollary that links layer-wise contribution to generalization performance under distribution shift.
@@
 Suppressing these layers’ contribution via learned residual scaling improves out-of-distribution generalization, while amplifying them degrades it.
-We introduce a residual scaling mechanism with learnable scalar coefficient parameters (Algorithm 1), inspired by prior work on adaptive residual modulation (Liu et al., 2019a; Menghani et al., 2024).
+We introduce a residual scaling mechanism with learnable scalar coefficient parameters, inspired by prior work on adaptive residual modulation (Liu et al., 2019a; Menghani et al., 2024).
 Transformer architectures inherently employ residual connections to iteratively refine representations.
@@
 $\beta_{\ell}$ is a learnable scalar parameter associated with transformer layer $\ell$, which modulates the contribution of that layer’s residual output to the model’s forward pass.
-Algorithm 1 Residual Scaling: Probing the Contribution of Transformer Blocks to the Generalization–Memorization Trade-off.
-1:Pretrained Transformer with $L$ layers, dataset $\mathcal{D}$, learning rate $\eta$
-2:Initialize $\beta_{1},\dots,\beta_{L}\leftarrow 1.0$ (trainable); freeze other weights
-3:for each training step do
-4:  Sample batch $(x,y)\sim\mathcal{D}$
-5:  $z^{(0)}\leftarrow\text{Embedding}(x)$
-6:  for $\ell=1$ to $L$ do
-7:   $r^{(\ell)}\leftarrow\text{block}^{(\ell)}(z^{(\ell-1)})$
-8:   $z^{(\ell)}\leftarrow z^{(\ell-1)}+\beta_{\ell}\cdot r^{(\ell)}$
-9:  end for
-10:  Compute loss $\mathcal{L}(z^{(L)},y)$
-11:  Update $\beta_{1},\dots,\beta_{L}$ using gradient descent
-12:end for
 We freeze model weights and optimize only the residual scaling coefficient parameters $\beta_{\ell}$, which are initialized to 1.
@@
 (b) GPT-2 Small on CLUTRR
-Figure 6: Residual scaling coefficients $\beta_{\ell}$ across all transformer layers.
+Figure 6: Residual scaling coefficients $\beta_{\ell}$ across all layers.
 ID training emphasizes later layers, while OOD training has lower final layer weight.
@@
 In contrast, under out-of-distribution (OOD) settings, the learned coefficients systematically have relative lower contribution of the final layers compared to the ID settings, revealing that improved generalization is associated with downweighting late-stage residual signals.
-More results are in Appendix E.
+Additional results are in Appendix E.
 This pattern supports the interpretation that late-layer residuals become increasingly specialized in memorized patterns tied to the training distribution.
 Together, these findings provide empirical support for the generalization ridge hypothesis, demonstrating that information flow in transformers reflects a trade-off between generalizable signals and memorized, distribution-specific features.
+6 Extension to Multi-Token Senario
+We further evaluate information flow under multi-token outputs using the CNN/DailyMail summarization dataset (1.0.0, test) (See et al., 2017; Hermann et al., 2015).
+We evaluate two pretrained base models: Qwen-2.5-0.5B and LLaMA-3.1-8B, generating 50 tokens per sample.
+Kernel Extension for Variable-Length Sequences.
+To handle multi-token targets, we extend our matrix-based mutual information (MI) estimation to variable-length sequences.
+Let $Y_{i}=\{y_{i}^{(t)}\}_{t=1}^{T_{i}}$ denote the target sequence for sample $i$.
+We employ the following kernel to compute sample-to-sample similarity:
+$K(Y_{i},Y_{j})=\frac{1}{T_{i}T_{j}}\sum_{t=1}^{T_{i}}\sum_{t^{\prime}=1}^{T_{j}}\kappa\!\left(y_{i}^{(t)},y_{j}^{(t^{\prime})}\right),$
+where $\kappa(\cdot,\cdot)$ is a Gaussian kernel.
+This formulation aggregates all pairwise token similarities between two target sequences into a single scalar, capturing the overall alignment between two token distributions.
+We quantify how much information each layer carries about the gold answer by computing $I(Z_{\ell}^{(t)};Y)$ in Figure 7, where $Z_{\ell}$ denotes the hidden state at the last token position at decoding step $t$ and $Y$ represents the full target sequence.
+(a) LLaMA on CNN/Daily for 50 generation steps.
+(b) Qwen on CNN/Daily for 50 generation steps.
+Figure 7: $I(Z_{\ell}^{(t)};Y)$ across layers (x-axis) and generation steps (left color bar), the right y-axis shows mutual information values.
+Values are plotted every ten steps.
+Lines denote the mean across five independent runs, with shaded regions indicating $\pm 1$ standard deviation.
+Mutual information across generation steps.
+We observe several consistent patterns in the mutual information profiles across generation steps.
+First, the overall magnitude of $I(Z_{\ell}^{(t)};Y)$ increases as generation progresses: later decoding steps tend to exhibit higher mutual information than earlier ones.
+This reflects the accumulation of information in the hidden representations as the model conditions on an increasingly informative prefix composed of its previously generated tokens.
+As generation progresses, the hidden state progressively approach the target sequence distribution.
+Second, despite differences in magnitude, the layer-wise structure of the mutual information curves remains consistent across generation steps.
+In particular, $I(Z_{\ell}^{(t)};Y)$ peaks at early-to-intermediate layers and gradually declines toward later layers.
+The concentration of mutual information in early-to-intermediate layers reflects their role in forming task-relevant representations under autoregressive decoding.
+These layers integrate information from the input and the evolving context into relatively abstract features that are shared across decoding steps, such as semantic content and task-level structure.
+In contrast, representations in later layers are increasingly shaped by step-specific decoding requirements, including local token selection, memorized pretrained corpus-specific surface patterns and distribution sharpening.
+While these representations are critical for producing the next output token, later layers increasingly compress information into a lower-entropy, decision-focused subspace, leading to reduced mutual information with the target sequence.
+Figure 8: Mutual information comparing stopwords/punctuation against content words.
+Figure 8 compares $I(Z_{\ell};Y)$ when generating content words versus stopwords and punctuation.
+Both models exhibit similar layer-wise trends across the two token types, with curves largely overlapping throughout early and middle layers, suggesting that intermediate layers perform semantic processing regardless of output token type.
+A divergence emerges in the final layers: content words show higher $I(Z_{\ell};Y)$ than stopwords.
+This reflects differences in lexical selection complexity.
+Content words draw from a large vocabulary space and require precise semantic discrimination, demanding more target information.
+Stopwords follow predictable syntactic patterns determined by local grammatical constraints, requiring less target information for selection.
 7 Conclusion
@@
 This phenomenon reflects a fundamental trade-off between generalization and memorization as information flows deeper into the model.
-Residual scaling experiments further corroborate this interpretation, demonstrating the functional specialization of different layers.
-Taken together, these findings position InfoRidge as a comprehensive framework for diagnosing how language models internally manage information during natural language generation, while also revealing the structural mechanisms that govern the balance between generalization and memorization in transformer architectures.
-Ethics Statement.
-This work adheres to the Code of Ethics.
-Our experiments use only open-source models and publicly available datasets under their respective open licenses, with no involvement of human subjects or sensitive data.
-We identify no foreseeable ethical risks.
-Reproducibility Statement.
-We ensure reproducibility by providing experimental and implementation details in Section 4 and Appendices BC.
-Full results with statistical significance are in Appendix E, and anonymous source code is included as supplementary material.
+A similar pattern is observed in the multi-token setting.
+Attention analysis, capacity analysis and residual scaling experiments offer complementary causal evidence supporting this observation.
+Taken together, these findings position InfoRidge as a comprehensive framework for diagnosing how language models internally manage information during natural language generation, while also revealing the structural mechanisms that govern the balance between generalization and memorization.
+Limitation
+Our analysis considers a limited set of datasets and models.
+In addition, the kernel design for multi-token scenarios may not optimally capture the ground-truth distribution.
+Future work will extend analysis to broader model families and tasks and explore improved kernel formulations.
 Appendix A Mathematical Details for Matrix-Based Information Estimation and Theoretical Foundations
@@
 A more uniform spectrum (i.e., higher entropy) suggests more diversity in the representation space, while a sharply peaked spectrum (i.e., low entropy) indicates redundancy or compression.
+Polynomial Kernel.
+Given a set of normalized representations $U=\{u_{i}\}_{i=1}^{N}$, the polynomial kernel Gram matrix is defined as $(G^{\mathrm{poly}}_{U})_{ij}=\left(u_{i}^{\top}u_{j}+c_{0}\right)^{p},$ where $p$ is the polynomial degree and $c_{0}$ is a constant bias term.
+Laplacian Kernel.
+Given the same representation set $U=\{u_{i}\}_{i=1}^{N}$, the Laplacian kernel Gram matrix is defined as $(G^{\mathrm{lap}}_{U})_{ij}=\exp\!\left(-\gamma\,\|u_{i}-u_{j}\|_{1}\right),$ where $\gamma>0$ is a kernel coefficient and $\|\cdot\|_{1}$ denotes the $\ell_{1}$ distance.
 A.2 Mutual Information Estimation
@@
 A more concentrated spectrum in $G_{UV}$ relative to $G_{U}$ and $G_{V}$ implies stronger dependence and thus higher mutual information.
-A.3 Min Wasserstein Generalization Bound (He et al., 2024)
-We restate the generalization bound proposed by He et al. (He et al., 2024), which characterizes generalization in terms of the Wasserstein distance between internal representations.
-Suppose that the loss function $\tilde{\ell}:\mathcal{Y}\times\mathcal{Y}\to\mathbb{R}_{\geq 0}$ is $\rho_{0}$-Lipschitz, and the activation function $\phi_{\ell}:\mathbb{R}\to\mathbb{R}$ is $\rho_{\ell}$-Lipschitz for each $\ell=1,\ldots,L$.
-Then:
-$\mathrm{gen}(P_{W|\mathcal{D}_{n}}\!,\!P_{X,Y})\!\leq\!\min_{\ell}\frac{\rho_{0}}{n}\sum_{i=1}^{n}\!\mathbb{E}_{W}\!\left[\!\left(1\!\vee\!\prod_{j=\ell+1}^{L}\!\rho_{j}\|W_{j}\|_{\mathrm{op}}\!\right)\!\cdot\!\mathcal{W}_{1}\!\left(\!P_{T_{\ell},i,Y_{i}|W}(\cdot|W),\!\ P_{T_{\ell},Y|W}(\cdot|W)\!\right)\!\right]$ (3)
-This result shows that generalization can be tightly controlled by the Wasserstein distance (Villani et al., 2008) between representations at a specific layer—referred to the generalization funnel layer.
-Connection to Our Work.
-The Min Wasserstein Generalization Bound (He et al., 2024) provides a theoretical foundation for our study by characterizing generalization in terms of distributional alignment at an intermediate layer.
-It motivates our analysis of information flow by suggesting that information is peaked at a specific layer—the generalization funnel.
-Our InfoRidge builds on this insight by quantifying predictive information across layers, and reveals that a specific intermediate layer exhibits peak mutual information and correlates with better generalization performance.
 Appendix B Dataset Overview and Statistics
-To evaluate information flow and generalization dynamics across model layers, we conduct experiments on three datasets with varying levels of complexity and structure: CLUTRR, ECQA, and a custom-designed Synthetic Arithmetic dataset.
-Table B.1 summarizes key dataset statistics.
-B.1 Dataset Overview
+Data Statistics
+We conduct experiments on four datasets with varying levels of complexity and structure: CLUTRR (Sinha et al., 2019), ECQA (Aggarwal et al., 2021), CNN/DailyMail (See et al., 2017; Hermann et al., 2015) and a custom-designed Synthetic Arithmetic dataset.
+Table 3 summarizes key dataset statistics.
 Table 3: Dataset Statistics
@@
 ECQA 7,598 21 1,090 21 2,194 21
+CNN/DailyMail - - - - 11,490 3,967
 Synthetic Arithmetic 10,000 9 2,000 9 2,000 9
-B.2 CLUTRR
+B.1 CLUTRR
 CLUTRR (Compositional Language Understanding and Text-based Relational Reasoning) (Sinha et al., 2019) is a diagnostic benchmark for evaluating relational inference in language models.
@@
 We use the task split “gen_train23_test2to10”, where the model is trained on clause lengths 2 and 3 and evaluated on lengths 2 through 10.
-B.3 ECQA
+B.2 ECQA
 ECQA (Explanations for CommonsenseQA) (Aggarwal et al., 2021) is a commonsense multiple-choice question-answering dataset, where each question is accompanied by 5 answer options.
+B.3 CNN/DailyMail
+CNN/DailyMail (See et al., 2017; Hermann et al., 2015) is an summarization dataset containing news articles as written by journalists at CNN and the Daily Mail.
 B.4 Synthetic Arithmetic Dataset
@@
 C.1 Prompt Construction
-All tasks are cast into a next-token generation format. The model receives a prompt and generate the next token. Below are construction strategies and examples for each dataset:
+All tasks are cast into a natural language generation format. The model receives a prompt and generate the next token. Below are construction strategies and examples for each dataset:
 CLUTRR
@@
 Target: B
+CNN / DailyMail
+[…]
+Prompt:
+(CNN)The Palestinian Authority officially became the 123rd member of the International Criminal Court on …
+The following is a one-paragraph highlight summary of the article:
+Target: Membership gives the ICC jurisdiction over alleged crimes committed in Palestinian territories since last June .
+Israel and the United States opposed the move, which could open the door to war crimes investigations against Israelis .
 Synthetic Arithmetic
@@
 Training converged in all settings.
-C.3 Information Estimation
-To estimate mutual information, we subsample between 50 and 200 test examples depending on model and task to achieve a stable result.
-C.4 Residual Scaling with Learnable $\beta_{\ell}$ Parameters
+C.3 Residual Scaling with Learnable $\beta_{\ell}$ Parameters
 We introduce a vector of learnable scalar weights $\beta=\{\beta_{1},\dots,\beta_{L}\}$ applied to residual connections in a frozen transformer:
@@
   • $\beta$ are initialized to 1 and updated using gradient descent without altering the architecture.
+The in-distribution (ID) setting is using the training split of “gen_train23_test2to10”.
+Out-of-distribution (OOD) is conducted on held-out CLUTRR configurations that do not overlap with the training distribution: for GPT models we evaluate on the test split of “gen_train23_test2to10” (lengths 2 through 10), while for Qwen and LLaMA we use the test split of “rob_train_sup_23_test_all_23”, which defines a more challenging OOD setting by jointly varying both the clause length and the status of noise facts introduced into the story, given the higher modeling capacity of these models.
+In all cases, tasks 1.2 and 1.3 are excluded from the OOD sets, as they overlap with the training configuration and are therefore treated as in-distribution rather than out-of-distribution.
 Training and Evaluation Settings
@@
 The ECQA (Aggarwal et al., 2021) dataset is also released under a CC BY 4.0 license by its authors as part of EMNLP 2021.
+CNN/DailyMail is under Apache-2.0 License.
 The Synthetic Arithmetic dataset is custom-designed by the authors and does not rely on any external or licensed data sources.
@@
 Appendix E Full Quantitative Results with Confidence Estimates
-To provide a comprehensive view of model behavior, we report the full experimental results for all models and datasets.
-This includes predictive information trends ($I(Z;Y)$), incremental information gain ($I(\Delta Z;Y)$) and residual scaling effects ($\beta_{\ell}$), with statistical significance reporting.
-E.1 Predictive Information
-To estimate mutual information between hidden representations and target labels, we compute matrix-based mutual information following the methodology in Appendix A.
-We subsample between 50 and 100 test examples per experiment and report results with statistical significance based on multiple random seeds.
+Predictive Information
+We compute mutual information following the methodology in Appendix A. We subsample between 50 and 100 test examples per experiment and report results with statistical significance based on 5 seeds.
 (a) GPT-2 Small on CLUTRR
@@
 (j) LLaMA-3.1-8B on Synthetic Arithmetic
-Figure 7: Predictive information $I(Z;Y)$ across different models and datasets exhibits an information peak, indicating a generalization ridge.
-In cases where the task is too simple relative to model capacity—such as the synthetic arithmetic task with LLaMA—this trend reflects an overfitting regime.
+Figure 9: Predictive information $I(Z;Y)$ across models and datasets.
 Lighter line colors represent later training epochs.
 Each curve shows the mean across 5 random seeds (0, 1, 2, 3, 42), and the shaded region denotes a 2-sigma (~96%) confidence interval.
-E.2 Incremental Information Gain
+In cases where the task is too simple relative to model capacity—such as the synthetic arithmetic task with LLaMA—this trend reflects an overfitting regime.
+Incremental Information Gain
 In addition to predictive information $I(Z^{(\ell)};Y)$, we compute the incremental information gain $I(\Delta Z^{(\ell)};Y)$ at each layer.
@@
 Figure 10: Incremental information gain $I(\Delta Z;Y)$ across different models and datasets with ~96% CI error bars.
-Across all models, we observe that the largest incremental information gain consistently occurs in intermediate layers—further supporting the emergence of a generalization ridge.
-E.3 Residual Scaling
-We introduce a residual scaling mechanism with learnable scalar coefficient parameters, inspired by prior work on adaptive residual modulation (Liu et al., 2019a; Menghani et al., 2024).
-Similar ideas of modulating internal computation have also been explored in parameter-efficient fine-tuning (PEFT) (Houlsby et al., 2019), Representation-Efficient Fine-Tuning (REFT) (Wu et al., 2024) and interpretability-driven control (Huang et al., 2024; Deng et al., 2025; Meng et al., 2022; Wu et al., 2023).
+Kernel Ablation
+To assess robustness with respect to kernel selection, we additionally test the Laplacian and Polynomial kernels (Figure 11).
+(a) Gaussian Kernel: Qwen-2.5-0.5B on ECQA.
+(b) Polynomial Kernel: Qwen-2.5-0.5B on ECQA.
+(c) Laplacian Kernel: Qwen-2.5-0.5B on ECQA.
+Figure 11: Different kernel shows similar pattern.
+Residual Scaling
 We present the complete set of residual scaling results, detailing the learned $\beta_{\ell}$ values across all transformer layers.
-These values reflect the relative contribution of each layer after optimizing the residual scaling coefficients while keeping all other model parameters frozen.
-$\beta_{\ell}$ are trained on in-distribution (ID) and out-of-distribution (OOD) data.
-We observe that in the ID setting, later layers tend to receive higher weights, consistent with memorization behavior.
-In contrast, OOD training consistently has lower final layers’ weight compared to the ID settings and shifts importance toward the intermediate layers.
-When the task is simple relative to the model’s capability, such as LLaMA on the synthetic task, the final-layer residual coefficients are similar for ID and OOD, consistent with Figure 7.
-The in-distribution (ID) setting is using the training split of “gen_train23_test2to10”.
-Out-of-distribution (OOD) is conducted on held-out CLUTRR configurations that do not overlap with the training distribution: for GPT models we evaluate on the test split of “gen_train23_test2to10” (lengths 2 through 10), while for Qwen and LLaMA we use the test split of “rob_train_sup_23_test_all_23”, which defines a more challenging OOD setting by jointly varying both the clause length and the status of noise facts introduced into the story, given the higher modeling capacity of these models.
-In all cases, tasks 1.2 and 1.3 are excluded from the OOD sets, as they overlap with the training configuration and are therefore treated as in-distribution rather than out-of-distribution.
 (a) GPT-2 Small on CLUTRR
@@
 Figure 12: Residual scaling coefficients $\beta_{\ell}$ across all transformer layers.
+Each curve shows the mean across five random seeds, and the shaded region denotes 1-sigma error bar.
 ID training emphasizes later layers, while OOD training has lower final layer weight.
-Each curve shows the mean across five random seeds (0, 1, 2, 3, 42), and the shaded region denotes 1-sigma error bar.
-Appendix F Attention Dynamics Across Layers Case Study
-To verify whether generalization ridge layer indeed correspond to semantically meaningful processing, we visualize the attention patterns as a more interpretable signal of where the model focuses.
-As shown in Figure 10, we visualize attention maps across layers.
-Figure 10: Attention map across layers (GPT-2 Small, Synthetic). At the generalization ridge layers, attention becomes more targeted, focusing on predictive tokens.
-In early layers (e.g., Layer 1), attention is diffuse and biased toward final tokens—reflecting reliance on position rather than true predict signals.
-By generalization ridge layers (e.g., Layer 11), attention becomes more targeted, concentrating on predictive tokens.
-This shift marks a transition from positional attention to semantically meaningful focus, suggesting that intermediate layers are increasingly capable of isolating task-relevant information from irrelevant content.
-In the final layers, attention regresses toward final tokens.
-This reversion is aligned with the observed decline in $I(Z_{\ell};Y)$.
-The resurgence of attention towards terminal tokens indicates a potential memorization scenario, where the model re-engages superficial positional strategies, possibly memorizing noise rather than further refining the generalized extraction of predictive signals.
-This attention trajectory further supports the generalization ridge hypothesis, highlighting a trade-off between generalization and memorization in the model’s representational strategy.
-Appendix G Incremental Information Gain Case Study
-To further illustrate how intermediate layers contribute to generalization, we analyze the semantic content introduced by residual transitions at different depths.
-Specifically, we decode the residual transition $\delta z$ at each layer using the language modeling (LM) head, projecting the incremental representation back into token space.
-This analysis allows us to inspect the linguistic shift introduced by each transformer block in isolation, and to assess whether the changes correspond to task-relevant predictions or superficial noises.
-G.1 Methodology
-For a fixed input, we compute residual transitions $\delta z^{(\ell)}$ at each layer and pass them through the model’s final linear projection (LM head) followed by a softmax.
-We record the top predicted tokens and their probabilities, and it reflects the directional change applied by layer $\ell$.
-G.2 Observations
-We decode residual transitions across layers and report the top shifted tokens by projecting $\Delta Z^{(\ell)}$ through the LM head.
-These shifts provide insight into how each layer modifies the model’s internal prediction trajectory.
-Layer 9 → 10: The top shifted tokens include ĠNine, Ġ10, Ġ4, Ġ8, and 8, which are all numerically aligned with the target prediction space in the synthetic arithmetic task.
-This indicates that the model is beginning to refine task-relevant numerical features at this depth.
-Layer 10 → 11: The shifted tokens become partially diluted, featuring punctuation and less informative symbols such as ,, Ġ, and ., alongside occasional task-relevant entries like Ġ12 and Ġ4.
-This indicates a transitional phase where the model continues to refine meaningful task-relevant features, yet begins to exhibit noise from frequent but semantically uninformative, such as punctuation, tokens.
-Layer 11 → 12: At the final layer transition, the top 5 shifted tokens become largely uninterpretable, including Ġcanaver, soDeliveryDate, enegger, 76561, and ILA.
-Conversely, the most negatively shifted tokens—Ġ4, Ġ3, Ġ5, Ġ6, Ġ1—correspond to plausible numerical predictions that were actively suppressed.
-This supports the hypothesis that final layers may overwrite generalizable abstractions with memorized or noise signals.
-These patterns are consistent with our mutual information analysis, which identifies intermediate layers as better semantically aligned with the prediction target—forming a ridge of generalization.
-G.3 Implication
-These decoding results reinforce our interpretation of the generalization ridge: intermediate layers contribute the most semantically informative updates to the model’s representation.
-The residual transitions thus serve as a useful lens for understanding how and where semantic meaning is introduced during forward propagation.
-Appendix H Activation Patching Experiment
-To assess which residual blocks causally support OOD prediction, we conduct a targeted activation-patching experiment.
-Let the model contain $L$ residual blocks with hidden states $Z_{\ell}\in\mathbb{R}^{d}$, and let $t^{*}$ denote the final context token (the last token before next-token prediction).
-For every OOD sequence $x_{\mathrm{ood}}$, we pair it with an ID sequence $x_{\mathrm{id}}$ sharing the same token structure.
-For each depth $\ell\in\{1,\dots,L\}$, we compute two forward passes: a standard OOD pass and a patched pass in which we replace only the hidden state of token $t^{*}$ at depth $\ell$ according to $Z_{\ell}(x_{\mathrm{ood}},t^{*})\leftarrow Z_{\ell}(x_{\mathrm{id}},t^{*}).$ All other activations remain unchanged.
-Let $logit_{\mathrm{base}}$ denote the logit assigned to the correct next token under the baseline OOD pass, and let $logit_{\mathrm{patch}}^{(\ell)}$ denote the corresponding logit under layer-$\ell$ patching.
-We measure causal impact using the logit drop $\Delta_{\ell}=logit_{\mathrm{base}}-logit_{\mathrm{patch}}^{(\ell)},$ where a positive value indicates that depth $\ell$ provides information necessary for correct OOD prediction.
-Table 5 reports the averaged logit drop for all $\ell=1,\dots,12$ (GPT-2 Small on Synthetic Dataset).
-The effects are negligible for layers 1–8, begin rising around layer 9, increase sharply at layer 10, peak at layer 11, and decline again at layer 12.
-This rise–peak–fall pattern closely matches the ridge structure observed in Figure 2, where $I(Z_{\ell};Y)$ increases through mid-to-late depth, reaches a maximum in the ridge region, and then decreases.
-The activation-patching experiment therefore provides supporting causal evidence that the ridge layers identified by IndoRidge are necessary for successful OOD generalization.
-Table 5: Average logit drop $\Delta_{\ell}$ (baseline logit - patched logit). Larger values indicate greater causal contribution to OOD prediction.
-Layer $\ell$ 1 2 3 4 5 6 7 8 9 10 11 12
-$\Delta_{\ell}$ -0.1037 -0.1039 -0.0327 -0.5405 -5.1663 -8.3992 -5.8914 -15.0996 -3.0998 9.3148 17.2022 15.5943
-Appendix I Extending Predictive Information to Multi-Token Outputs
-To complement our single-token analysis, we further evaluate information flow under multi-token outputs using the CNN/Daily Mail summarization dataset (1.0.0, test) (See et al., 2017; Hermann et al., 2015).
-Autoregressive next-token information.
-We follow the model’s natural decoding process.
-We analyze $I(Z_{\ell}^{(t)};\hat{Y}^{(t+1)})$ since there is no deterministic gold answer for each intermediate step in NLG, where for each generation step $t$, $Z_{\ell}^{(t)}$ is the hidden state of current token at layer $\ell$ and $\hat{Y}^{(t+1)}$ is the embedding of model’s predicted token at step $t+1$.
-Across steps, the pattern is consistent (Figure 11): information rises from early layers, peaks in the middle layers, and then declines toward top layers.
-This gives a clear information ridge, indicating that intermediate layers encode the most stable and task-relevant core information that supports the model’s decision-making (generalization).
-In particular, high $I(Z_{\ell}^{(t)};\hat{Y}^{(t+1)})$ at these layers means the model’s predictive direction is most strongly determined by that layer: layers whose representations consistently constrain the next-token prediction across many contexts must encode the shared, invariant structure that the model relies on to generalize.
-In contrast, upper layers exhibit behavior that might be related to pretrained surface patterns from corpus such as token co-occurrence, causing the representation to drift away from this decision-relevant structure (memorization).
-The first step is a notable outlier: it is the only step that conditions solely on the input prompt, and shows the sharpest and most pronounced ridge.
-Generating the first token requires a global transition from the prompt into the answer trajectory, which heavily engages the abstract, decision-relevant subspace in the middle layers.
-Crucially, this is also the point where information compression is most extreme, the model must condense all prompt-level semantics into a single initial token decision, therefore forms a substantial ridge.
-Once generation begins, later generation steps move into a more regular autoregressive regime, where each hidden state already contains a compressed summary of all previously generated content, and the continuation depends more on local context than on global restructuring.
-Consequently, the depth profiles become flatter and more stable across steps, although the mid-layer ridge remains visible.
-(a) Qwen-2.5-0.5B on CNN/Daily Mail for 50 generation steps.
-(b) LLaMA-3.1-8B on CNN/Daily Mail for 50 generation steps.
-Figure 11: $I(Z_{\ell}^{(t)};\hat{Y}^{(t+1)})$ across layers (x-axis) and generation steps (y-axis), where the left y-axis shows first step mutual information (bars) and the right y-axis shows mutual information for the rest steps.
-Pooled-answer information.
-We quantify how much information each layer carries about the gold answer by computing $I(Z_{\ell}^{(t)};Y^{\text{pool}})$, where $Y^{\text{pool}}$ is a fixed target vector obtained by average-pooling the embedding-layer representations of all ground-truth answer tokens.
-Across models and steps (Figure 12), a consistent pattern emerges: intermediate layers are where the hidden state is most strongly aligned with the gold-answer.
-(a) Qwen-2.5-0.5B on CNN/Daily Mail for 50 generation steps.
-(b) LLaMA-3.1-8B on CNN/Daily Mail for 50 generation steps.
-Figure 12: $I(Z_{\ell}^{(t)};Y^{\text{pool}})$ across layers (x-axis) and generation steps (y-axis), where the left y-axis shows first step mutual information (bars) and the right y-axis shows mutual information for the rest steps.
-Taken together, the two constructions reveal a coherent picture of depth specialization in multi-token generation.
-Intermediate layers are simultaneously the most informative about the model’s next-token decisions and the most aligned with the global answer.
-Thus, the two observed ridges are complementary projections of a single underlying phenomenon: intermediate layers concentrated most meaningful and decision-relevant information.
-In contrast, upper layers increasingly reflect pretrained corpus-specific surface patterns, causing the representation to drift away from this decision-relevant structure.
-(a) $I(Z_{\ell}^{(t)};\hat{Y}^{(t+1)})$: Qwen-2.5-0.5B on CNN/Daily Mail for 50 generation steps.
-(b) $I(Z_{\ell}^{(t)};\hat{Y}^{(t+1)})$: LLaMA-3.1-8B on CNN/Daily Mail for 50 generation steps.
-Figure 13: Mutual information comparing stopwords/punctuation against content words, showing that both stopwords/punctuation have consistently higher MI.
-We also plot the mutual information comparing stopwords/punctuation against content words, showing that both stopwords/punctuation have consistently higher MI (Figure 13).
-These tokens are more predictable because they follow strict grammatical and structural patterns, whereas content words depend more on semantic context and therefore exhibit greater variability.
-Appendix J Non-finetuned scenario
-Beyond fine-tuning scenarios, we also computed predictive-information $I(Z_{\ell};Y)$ curves in a non-training zero-shot regime and observed the same mid-layer information peak pattern, see Figure 14.
-Figure 14: GSM8K on Qwen2.5-0.5B shows an intermediate-layer MI rise and late-layer decline.
-Appendix K Kernel Ablation
-The matrix-based MI estimator requires a positive-definite Gram matrix, for which the Gaussian kernel is the standard choice.
-To assess robustness with respect to kernel selection, we additionally test the Laplacian and Polynomial kernels.
-While these kernels introduce quantitative shifts in the magnitude, the trend—including the rise–peak–decline structure and the location of the ridge—remains unchanged (Figure 15).
-Polynomial Kernel.
-Given a set of normalized representations $U=\{u_{i}\}_{i=1}^{N}$, the polynomial kernel Gram matrix is defined as
-$(G^{\mathrm{poly}}_{U})_{ij}=\left(u_{i}^{\top}u_{j}+c_{0}\right)^{p},$ (4)
-where $p$ is the polynomial degree and $c_{0}$ is a constant bias term.
-Laplacian Kernel.
-Given the same representation set $U=\{u_{i}\}_{i=1}^{N}$, the Laplacian kernel Gram matrix is defined as
-$(G^{\mathrm{lap}}_{U})_{ij}=\exp\!\left(-\gamma\,\|u_{i}-u_{j}\|_{1}\right),$ (5)
-where $\gamma>0$ is a kernel coefficient and $\|\cdot\|_{1}$ denotes the $\ell_{1}$ distance.
-(a) Gaussian Kernel: Qwen-2.5-0.5B on ECQA.
-(b) Polynomial Kernel: Qwen-2.5-0.5B on ECQA.
-(c) Laplacian Kernel: Qwen-2.5-0.5B on ECQA.
-Figure 15: Different kernel shows similar pattern.
-Appendix L Use of Large Language Models.
-Large language models (LLMs) were used solely as assistive tools for proofreading and improving clarity of writing.
+When the task is simple relative to the model’s capability, such as LLaMA on the synthetic task, the final-layer residual coefficients are similar for ID and OOD, consistent with Figure 9.

### v4

@@
 (d) LLaMA-3.1-8B on ECQA
-Figure 2: Evolution of predictive information $I(Z_{\ell};Y)$, with lighter curves indicating later epochs.
+Figure 2: Evolution of predictive information $I(Z_{\ell};Y)$, with lighter curves indicating later training steps.
 Each curve exhibits a three-phase trend: early layers rise, mid layers peak, and late layers decline.
@@
 $I(Z_{\ell};Y)$ continues to rise through the mid-to-upper layer, typically peaking before the final few blocks.
-For the GPT-2 Small on Synthetic Arithmetic dataset, the peak reaches $I\!\approx\!0.33$ in layers 10-11, coinciding with a jump from $\approx 0\%$ to $73\%$ ID accuracy and $57\%$ OOD accuracy.
+For the GPT-2 Small on Synthetic Arithmetic dataset, the peak reaches $I\!\approx\!0.9$ in layers 10-11, coinciding with a jump from $\approx 0\%$ to $85\%$ ID accuracy and $59\%$ OOD accuracy.
 These layers appear to play a critical role in synthesizing abstract features that are essential for generalization.
@@
 Out-Dist.
-Layer 1 0.0782 0.00 0.00 0.00 Layer 2 0.0776 0.00 0.00 0.00 Layer 3 0.0907 0.00 0.00 0.00 Layer 4 0.1156 0.00 0.00 0.00 Layer 5 0.1222 0.00 0.00 0.00 Layer 6 0.1275 0.00 0.00 0.00 Layer 7 0.1552 0.75 0.00 1.50 Layer 8 0.2594 19.45 0.00 40.70 Layer 9 0.2937 20.60 0.60 42.00 Layer 10 0.3295 64.25 72.90 56.90 Layer 11 0.3296 71.65 100.00 42.90 Layer 12 0.1501 70.45 100.00 41.40
+Layer 1 0.0585 0.00 0.00 0.00 Layer 2 0.0673 0.00 0.00 0.00 Layer 3 0.0969 0.00 0.00 0.00 Layer 4 0.1880 0.00 0.00 0.00 Layer 5 0.3038 0.00 0.00 0.00 Layer 6 0.3991 0.00 0.00 0.00 Layer 7 0.5497 4.00 0.10 7.90 Layer 8 0.8067 19.30 0.00 38.60 Layer 9 0.8604 22.35 1.30 43.40 Layer 10 0.9044 57.00 54.30 59.70 Layer 11 0.8650 70.35 85.10 55.60 Layer 12 0.1783 71.95 93.80 50.10
 Representational Compression (final layers).
@@
 Generalization Ridge Emerges Beyond a Depth Threshold
-(a) 8 layers: Monotonic Increase
+(a) 7 layers: Monotonic Increase
 (b) 9 layers: Shallow Peak
 (c) Full model: Gen Ridge
-Figure 3: Truncating GPT–2 to 8 layers removes the MI peak; 9–layer variants begin to exhibit a shallow peak, and the full 12–layer model shows a pronounced decline.
+Figure 3: Truncating GPT–2 to 7 layers removes the MI peak; 9–layer variants begin to exhibit a shallow peak, and the full 12–layer model shows a pronounced decline.
 We empirically vary the depth of GPT-2 Small models fine-tuned on the CLUTRR dataset to examine how model capacity shapes information dynamics, shown in Figure 3.
-When the model is truncated to 8 layers, $I(Z;Y)$ increases monotonically—indicating insufficient capacity to develop an information peak.
+When the model is truncated to 7 layers, $I(Z;Y)$ increases monotonically—indicating insufficient capacity to develop an information peak.
 Upon increasing the depth to 9 layers, a shallow peak emerges, signifying the threshold at which the model begins to distinguish generalized features from memorized signals.
@@
 (2) Final layers show reduced attention to signal tokens, suggesting a shift toward memorization to specific data point rather than predictive information abstraction.
-For example, in Qwen-2.5-0.5B on ECQA, signal attention peaks at Layer 17 (0.2458) but drops to 0.0066 in the final layer, where last-token attention dominates (0.1697).
+For example, in Qwen-2.5-0.5B on ECQA, signal attention peaks at Layer 17 (0.2307) but drops to 0.0104 in the final layer, where last-token attention dominates (0.5015).
 Table 2: Average attention statistics: (1) average attention scores over all tokens, (2) average attention to signal tokens, (3) the maximum signal attention and its corresponding layer, (4) signal token attention in the final layer and (5) last token attention in the final layer.
 Model Dataset Avg. Attn (All) Avg. Attn (Signal) Layer w/ Highest Avg. Signal Final Avg. Signal Final Avg. Last
-GPT-2 Small Synthetic 0.0227 0.0483 10 (0.0758) 0.0573 0.0400
-GPT-2 Medium CLUTRR 0.0086 0.0155 19 (0.0453) 0.0147 0.0889
-Qwen-2.5-0.5B ECQA 0.0225 0.0257 17 (0.2458) 0.0066 0.1697
-LLaMA-3.1-8B ECQA 0.0220 0.0379 21 (0.1276) 0.0168 0.3904
+GPT-2 Small Synthetic 0.0222 0.0409 8 (0.0740) 0.0562 0.0348
+GPT-2 Medium CLUTRR 0.0086 0.0086 18 (0.0270) 0.0124 0.1649
+Qwen-2.5-0.5B ECQA 0.0225 0.0317 17 (0.2307) 0.0104 0.2169
+LLaMA-3.1-8B ECQA 0.0220 0.0394 21 (0.1595) 0.0094 0.5015
 *We remove the first token attention score to mitigate attention sink effects.
@@
 Each $\beta_{\ell}$ controls the strength of the residual contribution from layer $\ell$, enabling the model to adaptively emphasize or suppress specific blocks:
-$z^{(\ell)}=z^{(\ell-1)}+\beta_{\ell}\cdot\text{block}^{(\ell)}(z^{(\ell-1)}),\qquad\beta_{\ell}\in\mathbb{R}_{\geq 0}.$
+$z^{(\ell)}=z^{(\ell-1)}+\beta_{\ell}\cdot\text{block}^{(\ell)}(z^{(\ell-1)}),\qquad\beta_{\ell}\in\mathbb{R}.$
 Definition (Residual Scaling Coefficient).
@@
 (a) GPT-2 Small on Synthetic Arithmetic
-(b) GPT-2 Small on CLUTRR
+(b) Qwen on ECQA
 Figure 6: Residual scaling coefficients $\beta_{\ell}$ across all layers.
@@
 For evaluation, test sequences are generated using values of $K$ from the range $[5,26]$ excluding 13 to simulate a distribution shift.
-In the residual $\beta_{\ell}$ analysis, we use $K=13$ for in-distribution (ID) training and $K=17$ for out-of-distribution (OOD) training, allowing for a controlled comparison between generalization and memorization behavior.
 The model receives the first 9 tokens as input and is trained to predict the signal component of the 10th token.
@@
 This controlled format enables manipulation of distributional properties by varying the modulus $K$.
-C.2 Finetuning Settings
-We fine-tuned all layers end-to-end using AdamW (learning rate $5\times 10^{-6}$, weight decay $0.01$) with a linear schedule (warmup ratio $0.1$) and early stopping on validation loss.
-Training converged in all settings.
-C.3 Residual Scaling with Learnable $\beta_{\ell}$ Parameters
+C.2 Residual Scaling with Learnable $\beta_{\ell}$ Parameters
 We introduce a vector of learnable scalar weights $\beta=\{\beta_{1},\dots,\beta_{L}\}$ applied to residual connections in a frozen transformer:
@@
   • $\beta$ are initialized to 1 and updated using gradient descent without altering the architecture.
-The in-distribution (ID) setting is using the training split of “gen_train23_test2to10”.
-Out-of-distribution (OOD) is conducted on held-out CLUTRR configurations that do not overlap with the training distribution: for GPT models we evaluate on the test split of “gen_train23_test2to10” (lengths 2 through 10), while for Qwen and LLaMA we use the test split of “rob_train_sup_23_test_all_23”, which defines a more challenging OOD setting by jointly varying both the clause length and the status of noise facts introduced into the story, given the higher modeling capacity of these models.
-In all cases, tasks 1.2 and 1.3 are excluded from the OOD sets, as they overlap with the training configuration and are therefore treated as in-distribution rather than out-of-distribution.
-Training and Evaluation Settings
+For Synthetic Arithmetic, the in-distribution (ID) setting is using $K=13$, the out-of-distribution (OOD) setting is using $K=[5,26]$ excluding 13.
+For CLUTRR, the in-distribution (ID) setting is using the “1.2” and “1.3” split of “gen_train23_test2to10”.
+Out-of-distribution (OOD) is conducted on held-out CLUTRR configurations that do not overlap with the training distribution: the “1.4” to “1.10” split of “gen_train23_test2to10”.
+For ECQA, we use training set as the ID setting and CLUTRR as the OOD setting.
+C.3 Training and Evaluation Settings
 Table 4: Performance (%) of different models across datasets.
 Model CLUTRR Synthetic ECQA
-GPT-2 Small 30.98 71.45
-GPT-2 Medium 34.21 68.35
-Qwen-2.5-0.5B 31.94 83.20 57.06
-LLaMA-3.1-8B 57.68 94.00 77.39
-We perform residual $\beta_{\ell}$ training on top of a model that has already been fine-tuned on the target dataset, keeping all previously learned weights frozen and optimizing only the $\beta_{\ell}$ parameters using HuggingFace’s Trainer API.
-We sweep over batch sizes from 4 to 32 depending on GPU memory constraints, and tune the learning rate between $5\times 10^{-5}$ and $5\times 10^{-6}$.
-We train for 3–10 epochs based on model capacity and task complexity.
-Optimization is performed using the AdamW optimizer with a weight decay of 0.01.
+GPT-2 Small 56.57 71.95
+GPT-2 Medium 62.11 73.15
+Qwen-2.5-0.5B 58.61 88.80 65.45
+LLaMA-3.1-8B 72.62 89.90 80.49
+We finetune the model using HuggingFace’s Trainer API with batch sizes of 16, the learning rate of $\alpha=5\times 10^{-6}$, and AdamW optimizer with weight decay of 0.01.
+We perform residual $\beta_{\ell}$ training on top of a model that has already been fine-tuned on the target dataset, keeping all previously learned weights frozen and optimizing only the $\beta_{\ell}$ parameters.
+We use batch sizes of 4, the learning rate of 0.001, Adam optimizer and epochs of 10.
 Evaluation is conducted by comparing the next predicted token with the next ground-truth target token.
@@
 Predictive Information
-We compute mutual information following the methodology in Appendix A. We subsample between 50 and 100 test examples per experiment and report results with statistical significance based on 5 seeds.
+We subsample 100 test examples per experiment to compute mutual information and report results with statistical significance based on 5 seeds.
 (a) GPT-2 Small on CLUTRR
@@
 Figure 9: Predictive information $I(Z;Y)$ across models and datasets.
-Lighter line colors represent later training epochs.
+Lighter line colors represent later training steps.
 Each curve shows the mean across 5 random seeds (0, 1, 2, 3, 42), and the shaded region denotes a 2-sigma (~96%) confidence interval.
@@
 Incremental Information Gain
-In addition to predictive information $I(Z^{(\ell)};Y)$, we compute the incremental information gain $I(\Delta Z^{(\ell)};Y)$ at each layer.
-This quantifies the information contribution for each transformer block.
+In addition to predictive information $I(Z^{(\ell)};Y)$, we compute the incremental information gain $I(\Delta Z^{(\ell)};Y)$ for each transformer block.
 (a) GPT-2 Small on CLUTRR
@@
 We present the complete set of residual scaling results, detailing the learned $\beta_{\ell}$ values across all transformer layers.
-(a) GPT-2 Small on CLUTRR
-(b) GPT-2 Small on Synthetic Arithmetic
-(c) Qwen-2.5-0.5B on CLUTRR
-(d) Qwen-2.5-0.5B on Synthetic Arithmetic
-(e) LLaMA-3.1-8B on CLUTRR
-(f) LLaMA-3.1-8B on Synthetic Arithmetic
+(a) GPT-2 Small on Synthetic Arithmetic
+(b) GPT-2 Medium on CLUTRR
+(c) Qwen-2.5-0.5B on ECQA
+(d) LLaMA-3.1-8B on ECQA
 Figure 12: Residual scaling coefficients $\beta_{\ell}$ across all transformer layers.
@@
 ID training emphasizes later layers, while OOD training has lower final layer weight.
-When the task is simple relative to the model’s capability, such as LLaMA on the synthetic task, the final-layer residual coefficients are similar for ID and OOD, consistent with Figure 9.